### v2

# HyperMARL: Adaptive Hypernetworks for Multi-Agent RL

Abstract
Adaptability is critical in cooperative multi-agent reinforcement learning (MARL), where agents must learn specialised or homogeneous behaviours for diverse tasks. While parameter sharing methods are sample-efficient, they often encounter gradient interference among agents, limiting their behavioural diversity. Conversely, non-parameter sharing approaches enable specialisation, but are computationally demanding and sample-inefficient. To address these issues, we propose HyperMARL, a parameter sharing approach that uses hypernetworks to dynamically generate agent-specific actor and critic parameters, without altering the learning objective or requiring preset diversity levels. By decoupling observation- and agent-conditioned gradients, HyperMARL empirically reduces policy gradient variance and facilitates specialisation within FuPS, suggesting it can mitigate cross-agent interference. Across multiple MARL benchmarks involving up to twenty agents—and requiring homogeneous, heterogeneous, or mixed behaviours—HyperMARL consistently performs competitively with fully shared, non-parameter-sharing, and diversity-promoting baselines, all while preserving a behavioural diversity level comparable to non-parameter sharing. These findings establish hypernetworks as a versatile approach for MARL across diverse environments. [^1] All code will be made available soon.
Multi-Agent Reinforcement Learning, Parameter Sharing, Specialisation, Behavioural Diversity
1 Introduction
Figure 1: HyperMARL Policy Architecture. Common Agent-ID conditioned shared MARL policy (left) vs HyperMARL (right), which uses a hypernetwork to generate agent-specific policies and decouples observation- and agent-conditioned gradients.
Specialist and generalist behaviours are critical to collective intelligence. In biological and social systems, individual specialisation enhances performance by allowing agents to excel at specific tasks (Woolley et al., 2015; Smith et al., 2008; Surowiecki, 2004), while generalist (homogeneous) behaviours promote adaptability and efficiency (Kassen, 2002; Williams & O’Reilly III, 1998). In Multi-Agent Reinforcement Learning (MARL), tasks similarly vary and require adaptability (Li et al., 2021; Bettini et al., 2024; Albrecht et al., 2024): some demand specialised agent behaviours, others benefit from homogeneous strategies, and some may require a blend of both (mixed behaviours).
Thus optimal performance is tied to being able to represent the required behaviours. No parameter sharing (NoPS) (Lowe et al., 2017) enables specialisation by using separate policy and critic networks per agent, but introduces significant computational overhead and sample inefficiency (Christianos et al., 2021). Conversely, full parameter sharing (FuPS) (Tan, 1993; Gupta et al., 2017; Foerster et al., 2016) tackles these inefficiencies by training a single shared network, but struggles in scenarios requiring diverse or specialised behaviours (Kim & Sung, 2023; Fu et al., 2022; Li et al., 2021).
Balancing FuPS efficiency with the capacity for diverse behaviours therefore remains a central open problem in MARL. Prior works have explored intrinsic-rewards (Li et al., 2021; Jiang & Lu, 2021), role-based allocations (Wang et al., 2020a, b), specialised architectures (Kim & Sung, 2023; Li et al., 2024; Bettini et al., 2024), or sequential updates (Zhong et al., 2024). However, these methods often alter the learning objective, require prior knowledge of optimal diversity levels, involve delicate hyperparameter tuning, or necessitate maintaining agent-specific parameters and sequential updates (see Section 7).
These challenges lead us to ask: Can we design a shared MARL architecture that flexibly supports both specialised and homogeneous behaviours—without modifying the learning objective or relying on prior knowledge of the optimal diversity level? A key difficulty with FuPS (particularly for diverse behaviours) is the gradient interference among agents, whereby their updates negatively impact each other’s learning (Christianos et al., 2021; Zhong et al., 2024).
To address this, we propose HyperMARL, a novel parameter sharing method that leverages hypernetworks (Ha et al., 2016) to generate agent-specific actor and critic parameters conditioned on agent IDs or embeddings (Figure 1). Hypernetworks are neural networks that generate the weights of another network conditioned on an input context vector. Crucially, they have already proven effective at resolving conflicting gradients in multi-task RL (Navon et al., 2020) and mitigating catastrophic forgetting in continual learning (von Oswald et al., 2020). In HyperMARL, they adaptively generate specialised or homogeneous policies and decouple observation- from agent-conditioned gradients (Section 4.4), a mechanism that empirically reduces gradient variance and improves policy updates (Sarafian et al., 2021). We further show that HyperMARL attains lower policy gradient variance than FuPS and that this decoupling is critical for specialisation (Section 6), suggesting it helps mitigate cross-agent gradient interference in FuPS.
We validate HyperMARL on diverse MARL benchmarks—including Dispersion and Navigation from the Vectorized Multi-Agent Simulator (VMAS) (Bettini et al., 2022), Multi-Agent MuJoCo (MAMuJoCo) (Peng et al., 2021), and SMAX (Rutherford et al., 2024)—across environments with two to twenty agents that require homogeneous, heterogeneous, or mixed behaviours. HyperMARL consistently matches or outperforms NoPS, FuPS, and diversity-promoting methods such as DiCo (Bettini et al., 2024) and HAPPO (Zhong et al., 2024). Furthermore, HyperMARL achieves NoPS-level behavioural diversity while using a shared architecture – a challenge for conventional FuPS methods.
Our contributions are as follows:
  • We propose HyperMARL (Section 4), a method that leverages agent-conditioned hypernetworks to adaptively learn diverse, homogeneous or mixed behaviours, without altering the learning objective or requiring preset diversity levels.
  • We empirically evaluate HyperMARL (Section 5) across diverse MARL environments with up to 20 agents. HyperMARL consistently achieves competitive mean episode returns compared to NoPS, FuPS, and diversity-focused methods such as DiCo and HAPPO, while achieving NoPS-level behavioural diversity.
  • We demonstrate that HyperMARL empirically reduces policy gradient variance compared to FuPS and facilitates specialisation, suggesting the importance of gradient decoupling for mitigating cross-agent interference (Section 5.2; Section 6). Additionally, HyperMARL scales efficiently by maintaining a near-constant number of parameters as the number of agents increases (Section C.2).
2 Background
We formulate the fully cooperative multi-agent systems addressed in our work as a Dec-POMDP (Oliehoek & Amato, 2016). A Dec-POMDP is a tuple, $\langle\mathbb{I},\mathbb{S},\{\mathbb{A}^{i}\}_{i\in\mathbb{I}},R,\{\mathbb{O}^{i}\}_{i\in\mathbb{I}},O,T,\rho_{0},\gamma\rangle$, where $\mathbb{I}$ is the set of agents of size $n=|\mathbb{I}|$, $\mathbb{S}$ is the set of global states with an initial state distribution $\rho_{0}$, $\mathbb{A}^{i}$ is the action space for agent $i$ where $\mathbb{A}=\times_{i}\mathbb{A}^{i}$ is the joint action space, $R:\mathbb{S}\times\mathbb{A}\to\mathbb{R}$ is the shared reward function, $\mathbb{O}^{i}$ is the observation space for agent $i$ with the joint observation space $\mathbb{O}=\times_{i}\mathbb{O}^{i}$, $O:\mathbb{O}\times\mathbb{A}\times\mathbb{S}\to[0,1]$ is the probability of joint observation $\mathbf{o}\in\mathbb{O}$, i.e. $O(\mathbf{o},\mathbf{a},s)=\Pr(\mathbf{o_{t}}|s_{t},\mathbf{a_{t-1}})$, $T:\mathbb{S}\times\mathbb{A}\times\mathbb{S}\to[0,1]$ is the state transition function i.e. $T\left(s,\mathbf{a},s^{\prime}\right)=\Pr\left(s_{t+1}\mid s_{t},\mathbf{a}_{t}\right)$ and $\gamma$ is the discount factor.
In this setting, each agent $i$ receives a partial observation $o_{t}^{i}\in\mathbb{O}^{i}$. These observations are accumulated into an action-observation history $h_{t}^{i}=\left(o_{0}^{i},a_{0}^{i},\ldots,o_{t-1}^{i},a_{t-1}^{i},o_{t}^{i}\right)$. Each agent $i$ acts based on their decentralized policies $\pi^{i}(a^{i}|h^{i})$. The joint history and joint action are defined as follows $\mathbf{h_{t}}=(h_{t}^{1},\ldots,h_{t}^{n})$ and $\mathbf{a_{t}}=(a_{t}^{1},\ldots,a_{t}^{n})$. The goal is to learn an optimal joint policy $\boldsymbol{\pi^{*}}=(\pi^{1*},...,\pi^{n*})$ that maximizes the expected discounted return as follows [^2] We use simplified notion here omitting explicit dependence on state transitions and distributions for brevity.,
$\boldsymbol{\pi}^{*}=\arg\max_{\boldsymbol{\pi}}\mathbb{E}_{s_{0}\sim\rho_{0},\;\mathbf{h}\sim\boldsymbol{\pi}}\left[G(\mathbf{h})\right]$ (1)
where $G(\mathbf{h})=\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},\mathbf{a}_{t})$.
The action-value and state-value functions of the joint policy $\boldsymbol{\pi}$ can be defined as:
$V(\mathbf{h})=\mathbb{E}_{\boldsymbol{\pi}}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},\mathbf{a}_{t})\,\Big{|}\,\mathbf{h}\right],$ (2)
$Q(\mathbf{h},\mathbf{a})=\mathbb{E}_{\boldsymbol{\pi}}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},\mathbf{a}_{t})\,\Big{|}\,\mathbf{h},\,\mathbf{a}_{0}=\mathbf{a}\right].$ (3)
2.1 Specialised Policies and Environments
Specialisation plays a key role in MARL, yet remains under-defined, so we define specialised environments and specialised policies.
Definition 2.1.
An environment is specialised if:
  1. 1.
    Distinct Agent Policies. The optimal joint policy $\boldsymbol{\pi}^{*}$ consists of at least two distinct agent policies, i.e., $\exists i,j\in\mathbb{I}$ such that $\pi^{i}\neq\pi^{j}$.
  2. 2.
    Non-Interchangeability. Any permutation $\sigma$ of the policies in $\boldsymbol{\pi}^{*}$, denoted as $\boldsymbol{\pi}^{\sigma}$, results in a weakly lower expected return:
    $\mathbb{E}_{\mathbf{h}\sim\boldsymbol{\pi}^{\sigma}}[G(\mathbf{h})]\leq\mathbb{E}_{\mathbf{h}\sim\boldsymbol{\pi}^{*}}[G(\mathbf{h})],$
    with strict inequality if the joint policies are non-symmetric (i.e., swapping any individual policy degrades performance).
For example, consider a specialised environment such as a football game, agents adopt distinct roles — such as ”attackers” and ”defenders” — to optimise the team’s performance. Permuting these roles (i.e., exchanging their policies) would typically lead to suboptimal results. Here, agents develop specialised policies by learning distinct, complementary behaviours essential for an optimal joint policy. Heterogeneous agents — with different action or observation spaces — are naturally specialised, while homogeneous agents can specialise by learning distinct policies critical to the team’s success. We analyse such specialised environments in Sections 3.1 and 5.2.
3 Are Independent or Fully Shared Policies Sufficient?
In this section, we introduce two simple matrix games and show that even in these basic settings, both fully independent and fully parameter shared policies have inherent limitations, highlighting the need for more nuanced approaches.
3.1 Specialisation and Synchronisation Game
Specialisation Game
Player 2
A B
Player 1 A (0.5, 0.5) (1, 1)
B (1, 1) (0.5, 0.5)
Table 1: 2-player specialisation game. Nash Equilibria in blue.
The Specialisation Game, inspired by the XOR game (Fu et al., 2022) and VMAS’s Dispersion (Bettini et al., 2022), is a two-player matrix game where each agent selects one of two food items. Payoffs are $0.5$ for matching choices and $1$ for different choices (payoff matrix in Table 1).
This game has two symmetric pure Nash equilibria on the anti-diagonal of the payoff matrix and can be extended to $n$ agents, with payoffs of 1 for unique actions and $1/n$ for matching actions. This satisfies the conditions of a Specialised Environment (Definition 2.1).
Synchronisation Game This is the inverse of the Specialisation Game, where players aim to choose the same food item. Payoffs are 1 for matching choices and 0.5 for different choices, resulting in pure Nash equilibria along the diagonal. For $n$ agents, matching actions receive a payoff of $1$, while differing choices receive $1/n$.
3.2 Limitations of Fully Shared and Independent Policies
Table 2: Final evaluation rewards (mean $\pm$ std) for REINFORCE variants in Specialisation and Synchronisation Games with 2, 4, and 8 agents. Bold values are within 1 std of the maximum. Fully shared or independent policies do not solve both games.
Specialisation Game
Method 2 4 8
NoPS 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
FuPS 0.50 $\pm$ 0.00 0.25 $\pm$ 0.00 0.13 $\pm$ 0.00
FuPS+ID 1.00 $\pm$ 0.00 0.95 $\pm$ 0.10 0.53 $\pm$ 0.09
Synchronisation Game
Method 2 4 8
NoPS 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.13 $\pm$ 0.00
FuPS 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
FuPS+ID 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.30 $\pm$ 0.35
In Appendix B (Proof 14) we show that a fully shared policy (without conditioning on agent IDs) cannot learn the optimal joint behaviour in the 2-player Specialisation Game. To assess this limitation in practice, we compare three REINFORCE (Williams, 1992) variants in both games with $n=2,4,8$ agents: NoPS (No Parameter Sharing), FuPS (Fully Parameter Sharing), and FuPS+ID (FuPS with one-hot agent IDs). All policies use single-layer neural networks with controlled parameter counts (see Appendix F for details).
Table 2 shows that NoPS consistently achieves optimal performance in the Specialisation Game, while FuPS performance declines as the number of agents increases. Conversely, in the Synchronisation Game, FuPS performs optimally, whereas NoPS fails with $n=8$. FuPS+ID initially performs well in both games but deteriorates at larger scales, indicating that ID conditioning alone is insufficient for fully distinct or uniform behaviours.
Overall, these experiments demonstrate that neither fully shared nor fully independent policies provide a universal solution—even in simple matrix games. This highlights the need for more flexible MARL strategies that handle diverse task demands. In Section 5, we show similar findings for modern MARL algorithms and more sophisticated environments.
4 HyperMARL
In this section, we introduce HyperMARL, an approach that uses agent-conditioned hypernetworks to learn diverse or homogeneous policies end-to-end, without modifying the standard RL objective or requiring prior knowledge of the task’s optimal diversity. By operating within a fully shared paradigm, HyperMARL leverages shared gradient information while enabling specialisation through the decoupling of observation- and agent-conditioned gradients. We present the pseudocode in Section C.1.
4.1 Hypernetworks for MARL
Hypernetworks (Ha et al., 2016) are neural networks that generate the weights of another network conditioned on an input context vector. In HyperMARL, we use them to adaptively produce per-agent policy and critic parameters.
As illustrated in Figure 1, for any agent $i$ with context $e^{i}$ (i.e., either a one-hot encoded ID or a learned embedding), the hypernetworks generate the agent-specific parameters:
$\theta^{i}=h_{\psi}^{\pi}(e^{i}),\quad\phi^{i}=h_{\varphi}^{V}(e^{i}),$ (4)
where $h_{\psi}^{\pi}$ and $h_{\varphi}^{V}$ are the hypernetworks for the policy and critic, respectively. The parameters $\theta^{i}$ and $\phi^{i}$ define each agent’s policy $\pi_{\theta^{i}}$ and critic $V_{\phi^{i}}$, dynamically enabling either specialised or homogeneous behaviours as required by the task.
We explore two types of hypernetworks in HyperMARL: Linear and Multilayer Perceptron (MLP) Hypernetworks, balancing expressiveness and computational efficiency.
Linear Hypernetworks Linear hypernetworks are a straightforward instantiation of HyperMARL. Given a one-hot agent ID, $\mathds{1}^{i}\in\mathbb{R}^{1\times n}$, a linear hypernetwork $h_{\psi}^{\pi}$ generates agent-specific parameters $\theta^{i}$ as  [^3] For conciseness we only show the policy parameters in this section.:
$\theta^{i}=h_{\psi}^{\pi}(\mathds{1}^{i})=\mathds{1}^{i}\cdot W+b$ (5)
where $W\in\mathbb{R}^{n\times m}$ is the weight matrix (with $m$ the per-agent parameter dimensionality and $n$ is the number of agents), and $b\in\mathbb{R}^{1\times m}$ is the bias vector. Since $\mathds{1}^{i}$ is one-hot encoded, each $\theta^{i}$ corresponds to a specific row of $W$ plus the shared bias $b$. If there is no shared bias term, this effectively replicates training of separate policies for each task (in our case, for each agent) (Beck et al., 2023), since there are no shared parameters and gradient updates are independent.
MLP Hypernetworks for Expressiveness While linear hypernetworks can be effective (Beck et al., 2023; Chang et al., 2020), they may lack the capacity to capture complex relationships between agents’ roles and their corresponding weights. To enhance expressiveness, we incorporate hidden layers and non-linear activations (e.g., ReLU) into the hypernetworks as follows:
$\theta^{i}=h_{\psi}^{\pi}(e^{i})=f_{\psi_{1}}^{\pi}\bigl{(}g_{\psi_{2}}^{\pi}(e^{i})\bigr{)}$ (6)
where $e^{i}$ is the agent ID or embedding for agent $i$, $g_{\psi_{2}}^{\pi}$ is the multilayer perceptrons (MLPs) that produce intermediate representations with non-linear activations, and $f_{\psi_{1}}^{\pi}$ and is the final linear transformations that generate the policy parameters.
This additional expressiveness may be critical in domains where roles overlap or where nuanced specialisations emerge dynamically. Unlike linear hypernetworks with one-hot agent IDs, MLP hypernetworks do not guarantee distinct weights for each agent. Additionally, they increase the total number of trainable parameters, necessitating a careful balance between expressiveness and computational overhead.
4.2 Agent Embeddings
The agent embedding $e^{i}$ represents the input to the hypernetwork. For our linear hypernetworks, we use one-hot encoded agent IDs as embeddings, resulting in a mix of agent-independent weights and shared weights as discussed in Section 4.1. For our MLP hypernetworks, we use learned agent embeddings, initialised orthogonally. These embeddings are optimised end-to-end alongside the hypernetwork, allowing agents to learn distinct or similar embeddings based on whether homogeneous or specialised policies are required.
4.3 Hypernetwork Initialisation
HyperMARL initialises its hypernetworks to ensure that the initial agent-specific parameters $\theta^{i}$ and $\phi^{i}$ adhere to the same distribution as standard networks without hypernetworks. For example, PPO commonly uses orthogonal initialisation (Engstrom et al., 2019), so HyperMARL generates orthogonally initialised weights for each agent, preserving the same fan in/out as the conventional network. This alignment with standard deep RL initialisation schemes facilitates stable learning and can be applied using any initialisation strategy.
4.4 Gradient Decoupling in HyperMARL
A core difficulty in FuPS is cross-agent gradient interference (Christianos et al., 2021; Zhong et al., 2024). In HyperMARL, gradient decoupling helps mitigate this issue.
Consider a fully cooperative MARL setting with a centralised critic, we can formulate the policy gradient for agent $i$ as follows (Albrecht et al., 2024; Kuba et al., 2021):
$\nabla_{\theta^{i}}J(\theta^{i})\;=\;\mathbb{E}_{\mathbf{h_{t}},\mathbf{a_{t}}\sim\boldsymbol{\pi}}\Bigl{[}A(\mathbf{h_{t}},\mathbf{a_{t}})\,\nabla_{\theta^{i}}\log\pi_{\theta^{i}}(a_{t}^{i}\mid h_{t}^{i})\Bigr{]},$
where $\mathbf{h_{t}}$ and $\mathbf{a_{t}}$ are the joint histories and joint actions for all agents, $\theta^{i}$ denotes the parameters of agent $i$, and $A(\mathbf{h_{t}},\mathbf{a_{t}})=Q(\mathbf{h_{t}},\mathbf{a_{t}})-V(\mathbf{h_{t}})$ is the advantage function.
When using shared parameters $\theta$, this can reframed as:
$\nabla_{\theta}J(\theta)=\sum_{i=1}^{I}\mathbb{E}_{\mathbf{h_{t}},\mathbf{a_{t}}\sim\boldsymbol{\pi}}\left[A(\mathbf{h_{t}},\mathbf{a_{t}})\,\nabla_{\theta}\log\pi_{\theta}\left(a_{t}^{i}\mid h_{t}^{i},id\right)\right],$ (7)
where $id$ is our agent ID.
In HyperMARL, each agent’s policy parameters $\theta^{i}$ are generated by a hypernetwork $h_{\psi}^{\pi}$ conditioned on an agent embedding $e^{i}$:
$\theta^{i}\;=\;h_{\psi}^{\pi}\bigl{(}e^{i}\bigr{)}.$
Hence, the gradient with respect to $\psi$ becomes:
$\nabla_{\psi}J(\psi)\;=\;\sum_{i=1}^{I}\mathbb{E}_{\mathbf{h_{t}},\mathbf{a_{t}}\sim\boldsymbol{\pi}}\Bigl{[}A(\mathbf{h_{t}},\mathbf{a_{t}})\,\nabla_{\psi}\log\pi_{\theta^{i}}(a_{t}^{i}\mid h_{t}^{i})\Bigr{]}.$
Applying the chain rule:
$\nabla_{\psi}\,\log\pi_{\theta^{i}}(a_{t}^{i}\mid h_{t}^{i})\;=\;\underbrace{\nabla_{\theta^{i}}\,\log\pi_{\theta^{i}}(a_{t}^{i}\mid h_{t}^{i})}_{\text{Observation-Conditioned}}\;\cdot\;\underbrace{\nabla_{\psi}\,h_{\psi}^{\pi}(e^{i})}_{\text{Agent-Conditioned}},$
leading to
$\nabla_{\psi}J(\psi)$ $=\sum_{i=1}^{I}\mathbb{E}_{\mathbf{h_{t}},\mathbf{a_{t}}\sim\boldsymbol{\pi}}\Bigl{[}A(\mathbf{h_{t}},\mathbf{a_{t}})\,\nabla_{\theta^{i}}\log\pi_{\theta^{i}}(a_{t}^{i}\mid h_{t}^{i})$
$\quad\cdot\nabla_{\psi}h_{\psi}^{\pi}(e^{i})\Bigr{]}.$ (8)
This factorisation separates:
Observation-Conditioned Gradients ($\nabla_{\theta^{i}}\!\log\pi_{\theta^{i}}$), which reflect how each policy updates in response to local observations.
Agent-Conditioned Gradients ($\nabla_{\psi}\,h_{\psi}^{\pi}(e^{i})$), which reflect how the hypernetwork generates each agent’s parameters from each agent’s embedding $e^{i}$.
These gradients thus remain partially decoupled, reducing cross-agent interference. Related work in Meta-RL shows that similar decoupling lowers policy gradient variance and stabilises learning (Sarafian et al., 2021). In Section 5.2, we observe a comparable reduction in variance with HyperMARL. Additionally, ablation studies in Section 6 show that removing gradient decoupling significantly degrades performance, highlighting its essential role in HyperMARL’s effectiveness.
5 Experiments
We evaluate HyperMARL on two primary questions:
Q1: Specialised Policy Learning: Can HyperMARL effectively learn specialised policies (see Section 2.1) using a shared hypernetwork, without altering the learning objective or relying on preset diversity levels?
Q2: Effectiveness in Homogeneous Tasks: Is HyperMARL competitive in environments that necessitate homogeneous behaviours?
5.1 Experimental Setup
Table 3: Summary of MARL environments used to evaluate HyperMARL.
Environment Agents Action Space Behaviour
Dispersion 4 Discrete Heterogeneous
Navigation 2, 4, 8 Continuous Homogeneous, Heterogeneous, Mixed
MAMuJoCo 2–17 Continuous Heterogeneous
SMAX 2–20 Discrete Homogeneous
Environments We use diverse cooperative MARL tasks (Table 3), varying in complexity, agent count (2 to 20), and required behaviours (homogeneous, heterogeneous, or mixed). Full details are in Appendix D.1.
MARL Baselines We focus on PPO variants, given their success in MARL (De Witt et al., 2020; Yu et al., 2022). Comparisons include IPPO (De Witt et al., 2020), which learns an actor and critic conditioned on individual observations, and MAPPO (Yu et al., 2022), which trains a centralised critic conditioned on the global state or concatenated agent observations. We use NoPS and FuPS variants of these methods, with all FuPS variants being conditioned on one-hot encoded agent IDs.
(a) IPPO
(b) MAPPO
Figure 2: Performance of IPPO and MAPPO on Dispersion after 20 million timesteps. We show the Interquartile Mean (IQM) of the Mean Episode Return and the 95% Stratified Bootstrap Confidence Intervals (CI) using (Agarwal et al., 2021). Hypernetworks achieve comparable performance to NoPS, while FuPS struggle with specialisation. Sample efficiency plots are provided in Appendix E.1.1.
(a) Policy Diversity using SND with Jenson-Shannon distance. Hypernetworks match NoPS-level diversity while sharing parameters. Bold indicates highest SND, asterisks show overlaps in std.
Method Algorithm Mean ± Std Dev
IPPO MLP Hypernet $6.3\times 10^{-4}\pm 2.0\times 10^{-5}$
Linear Hypernet $6.4\times 10^{-4}\pm 7.0\times 10^{-6}$
NoPS $\mathbf{6.9\times 10^{-4}\pm 7.0\times 10^{-6}}$
FuPS $2.4\times 10^{-4}\pm 1.1\times 10^{-4}$
MAPPO MLP Hypernet $\mathbf{6.7\times 10^{-4}\pm 5.0\times 10^{-6}}$
Linear Hypernet $6.6\times 10^{-4}\pm 1.2\times 10^{-5*}$
NoPS $6.6\times 10^{-4}\pm 1.5\times 10^{-5*}$
FuPS $3.7\times 10^{-4}\pm 8.1\times 10^{-5}$
(b) Average policy gradient variance across IPPO (left) and MAPPO (right). Our hypernetworks have lower variance than FuPS.
Privileged Baselines Beyond NoPS, we include two privileged baselines designed for heterogeneous policies. Diversity Control (DiCo) (Bettini et al., 2024) partitions each policy into shared and agent-specific components and enforces a target diversity level (via Wasserstein distances). Heterogeneous-Agent PPO (HAPPO) (Zhong et al., 2024) uses a shared centralised critic and updates individual actors sequentially, incorporating prior actor updates to allow agent heterogeneity. In contrast, HyperMARL uses fully shared weights, parallel updates, and no preset diversity levels. Table 7 summarises baseline selection criteria. Implementation and hyperparameters follow authors’ defaults unless otherwise stated (see Appendix F), and HyperMARL uses the same observations and generates architectures with equivalent capacity to the baselines. Training and evaluation protocols are consistent across all experiments (Appendix D.1.1).
5.1.1 Measuring Policy Diversity
To measure the diversity of the policies we System Neural Diversity (SND) (Bettini et al., 2023) (Equation 9) with Jensen-Shannon distance, where 0 indicates identical policies and 1 indicates maximal diversity. Details in Appendix D.1.2.
5.2 Q1: Specialised Policy Learning
Learning Diverse Behaviour (Dispersion) Figures 2(a) and 2(b) show that parameter sharing variants (IPPO-FuPS, MAPPO-FuPS – ($\bullet$)) fail to learn the diverse policies required by Dispersion, while their non-parameter sharing counterparts (IPPO-NoPS, MAPPO-NoPS–($\bullet$)) converge to the optimal policy. These results corroborate that parameter sharing methods struggle in environments requiring agent specialisation. In contrast, our shared linear and MLP hypernetworks ($\bullet$, $\bullet$) match NoPS performance, suggesting that a shared hypernetwork can effectively support agent specialisation.
Policy Diversity SND measurements (Table 3(a)) show that IPPO-FuPS and MAPPO-FuPS exhibit lower behavioural diversity than their NoPS counterparts, aligning with their weaker performance results. Notably, both linear and MLP hypernetworks achieve diversity levels comparable to NoPS despite using a shared architecture. Linear hypernetworks with one-hot agent IDs enforce weight seperation, while MLP hypernetworks have to dynamically learn diverse policies without explicit separation of weights or gradients.
Gradient Variance To investigate HyperMARL’s effect on training stability, we compute the average policy gradient variance by calculating the variance across all actor parameters at each update and averaging it over training. As shown in Figure 3(b), hypernetworks in both IPPO and MAPPO exhibit lower mean policy gradient variance than FuPS. This aligns with their ability to learn diverse behaviours and may stem from decoupling observation- and agent-condition gradients (Section 4.4), similar to Meta-RL findings (Sarafian et al., 2021).
Table 4: Mean episode return in MAMuJoCo for MAPPO variants. IQM of the mean episode returns with 95% stratified bootstrap CI. Bold denotes the highest score per scenario; * indicates overlap with the highest score within one CI. HyperMARL (MLP variant) is competitive with methods that use independent actors.
Scenario HAPPO FuPS Ind. Actors HyperMARL
Humanoid-v2 17x1 6501.15* (3015.88, 7229.79) 566.12 (536.36, 603.01) 6188.46* (5006.13, 6851.74) 6544.10 (3868.00, 6664.89)
Walker2d-v2 2x3 4748.06* (4366.94, 6230.81) 4574.39* (4254.21, 5068.32) 4747.05* (3974.76, 6249.58) 5064.86 (4635.10, 5423.42)
HalfCheetah-v2 2x3 6752.40* (6130.42, 7172.98) 6771.21* (6424.94, 7228.65) 6650.31* (5714.68, 7229.61) 7063.72 (6696.30, 7325.36)
Ant-v2 4x2 6031.92* (5924.32, 6149.22) 6148.58 (5988.63, 6223.88) 6046.23* (5924.62, 6216.57) 5940.16* (5485.77, 6280.59)
(a) All Agents with Same Goals
(b) All Agents with Different Goals
(c) Mixed Goals (Half Agents Share Goals)
Figure 4: Average Reward in Navigation for IPPO Variants. IQM and 95% CI of mean reward for IPPO Baselines, HyperMARL (linear and MLP), and DiCo in the Navigation environment across goal configurations. HyperMARL maintains strong IQM performance, particularly with more agents. Sample efficiency plots in Appendix E.3.
Diversity at Complexity and Scale (MAMuJoCo) Dispersion is designed to test diverse behaviours (Bettini et al., 2022), but we further evaluate HyperMARL on the more challenging heterogeneous MAMuJoCo environment. Following Zhong et al. (2024), we use a centralised but non-shared critic, applying HyperMARL only to the actor.
Table 4 shows that HyperMARL (MLP variant) is competitive with all baselines across various MAMuJoCo scenarios. Unlike HAPPO and MAPPO (independent actors), HyperMARL uses a shared actor and parallel updates, and yet manages strong performance. This is particularly notable in Humanoid-v2 17x1(Figure 5), where 17 agents to coordinate in a Humanoid control task, a notoriously difficult heterogeneous task(Zhong et al., 2024). This demonstrates HyperMARL’s ability to handle large-scale specialisation with a shared architecture.
Figure 5: Performance on the 17-agent Humanoid task. IQM of the mean episode returns with 95% stratified bootstrap CI. HyperMARL is able to match the performance of methods that leverage independent actors, while using a shared actor in this challenging heterogeneous task.
Adaptability (Navigation) We evaluate HyperMARL against DiCo (Bettini et al., 2024) in the Navigation environment (Bettini et al., 2022). For two agents ($n=2$), we adopt DiCo’s optimal diversity levels and replicate its hyperparameters. For larger teams ($n>2$), we perform hyperparameter sweeps for diversity levels and learning parameters across all methods (see Tables 14 and 15). Although DiCo defines optimal diversity for two agents, identifying suitable levels for larger team sizes proved challenging.
Figure 4 presents the IQM of the mean reward across different goal setups. HyperMARL consistently demonstrates higher average returns across all configurations, outperforming DiCo and matching or exceeding NoPS and FuPS performance for most settings. Unlike in Dispersion, IPPO-FuPS is competitive with NoPS and HyperMARL in the Navigation environment for tasks requiring diverse behaviours (mixed or different goals – Figures 4(b) and 4(c)) with two and four agents. We hypothesise that this arises from the dense rewards in Navigation, in contrast to the sparse rewards in Dispersion. However, when scaling to eight agents, HyperMARL outperforms all baselines, demonstrating its effectiveness in handling more complex coordination challenges.
Figure 6: IQM and 95% CI of mean return in SMAX. Performance of FuPS Recurrent IPPO and MAPPO, HyperMARL (MLP) on SMAX. HyperMARL performs comparably to these baselines.
5.3 Q2: Effectiveness in Homogeneous Tasks
SMAX  Finally, we test HyperMARL (MLP variant) on SMAX, comparing against the FuPS baseline from JaxMARL. Prior work (Yu et al., 2022; Fu et al., 2022) has found FuPS optimal in such environments. Figure 6 shows that across simple (2s3z), hard (3s5z), and SMACv2-style maps (smacv2_10, smacv2_20), HyperMARL is comparable to the FuPS baseline. Alongside results in Navigation (Figure 4(a)), this indicates that HyperMARL is also effective in tasks requiring homogeneous behaviours, large observation spaces, and coordination among many agents, reinforcing its effectiveness across diverse MARL scenarios.
6 Ablations
To better understand the importance of the different components in HyperMARL, we conduct an ablation study on four key variations of HyperMARL (MAPPO-MLP) in the Humanoid-v2 17x1 (Section 5.2). First, HyperMARL-S reduces the hypernetwork hidden layer size from 64 to 32, investigating the impact of network capacity. Second, HyperMARL w/ One-Hot replaces learned agent embeddings with one-hot IDs (Section 4.2). Third, HyperMARL w/o RF omits the initialisation scaling described in Section 4.3. Finally, HyperMARL w/o GD removes gradient decoupling (Section  4.4) by conditioning the hypernetwork on both observations and agent embeddings.
Figure 7 shows that HyperMARL-S converges more slowly but ultimately reaches an IQM close to the baseline HyperMARL, while learned agent embeddings offer better final performance than one-hot IDs (HyperMARL w/ One-Hot). Crucially, removing either gradient decoupling (HyperMARL w/o GD) or initialisation scaling (HyperMARL w/o RF) severely impairs learning, confirming their essential roles in HyperMARL ability to learn specialised policies. We see similar findings in ablations in Dispersion (Section 18).
Figure 7: Ablation results comparing HyperMARL and its variants in Humanoid-v2 17x1. Gradient decoupling and initialization are crucial for maintaining performance.
7 Related Work
Hypernetworks in RL and MARL. Hypernetworks have proven effective in meta-learning, multi-task learning, and continual learning in single-agent settings (Beck et al., 2023, 2024; Sarafian et al., 2021; Huang et al., 2021). In MARL, QMIX (Rashid et al., 2020) used hypernetworks (conditioned on a global state) to mix per-agent Q-values; however, each agent’s own network remained a standard MLP or GRU. Parallel work, CASH (Fu et al., 2025), conditions hypernetworks on local observations and team capabilities for zero-shot generalization with heterogeneous action spaces. In contrast, we focus on agent-conditioned hypernetworks for adaptive specialisation within a fixed state-action setting, leveraging gradient decoupling (absent in CASH) that we find critical for specialised behaviours.
Variants of Parameter Sharing. Selective Parameter Sharing (SePS) (Christianos et al., 2021) shares weights between similar groups of agents, identified via trajectory clustering. Pruning methods (Kim & Sung, 2023; Li et al., 2024) split a single network into agent-specific subnetworks.
Learning Diverse Policies. Shared parameters often limit policy diversity (Christianos et al., 2021; Kim & Sung, 2023; Fu et al., 2022; Li et al., 2021). Proposed solutions include: (1) maximizing mutual information between agent IDs and trajectories (Li et al., 2021), (2) role-based methods (Wang et al., 2020a, b), or (3) methods that use structural modifications or constraints to induce diversity in agent policies (Kim & Sung, 2023; Bettini et al., 2024; Li et al., 2024). Outside FuPS/NoPS, HAPPO (Zhong et al., 2024) uses a non-shared centralised critic with individual actors updated sequentially to learn heterogeneous behaviours.
8 Conclusion and Discussion
We introduced HyperMARL, an approach that uses agent-conditioned hypernetworks to generate per-agent actor and critic parameters without modifying the standard RL objective or requiring preset diversity levels. Our results show it can adaptively learn specialised, homogeneous, or mixed behaviours in settings with up to 20 agents. We also observe a link between HyperMARL’s performance and reduced policy gradient variance, underscoring the importance of decoupling observation- and agent-conditioned gradients. Overall, these findings establish HyperMARL as a versatile and promising architecture for diverse MARL tasks. Future work could explore more parameter-efficient hypernetworks (e.g., chunked variants (von Oswald et al., 2020; Chauhan et al., 2024)) to enable large-scale applications.
Impact Statement
This paper presents work whose goal is to advance the field of Multi-Agent Reinforcement Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Measuring Behavioural Diversity
A.1 Quantifying Team Diversity
We quantify policy diversity using System Neural Diversity (SND) (Bettini et al., 2023), which measures behavioural diversity based on differences in policy outputs:
$\operatorname{SND}\left(\left\{\pi^{i}\right\}_{i\in\mathbb{I}}\right)=\frac{2}{n(n-1)|\mathcal{O}|}\sum_{i=1}^{n}\sum_{j=i+1}^{n}\sum_{o\in\mathcal{O}}D\left(\pi^{i}(o),\pi^{j}(o)\right).$ (9)
where $n$ is the number of agents, $\mathcal{O}$ is a set of observations typically collected via policy rollouts, $\pi^{i}(o_{t})$ and $\pi^{j}(o_{t})$ are the outputs of policies $i$ and $j$ for observation $o_{t}$, and $D$ is our distance function between two probability distributions.
In contrast to  (Bettini et al., 2023), we use Jensen-Shannon Distance (JSD) (Endres & Schindelin, 2003; Lin, 1991) as $D$, rather than the Wasserstein metric (Vaserstein, 1969). As shown in Appendix A.2, JSD is a robust metric for both continuous and discrete cases, and provides a more reliable measure of policy distance.
A.2 Finding a Suitable Distance Function for Policy Diversity
The choice of distance function $D$ in Equation 9 is crucial for accurately measuring policy diversity. In MARL, policies are often represented as probability distributions over actions, making the choice of distance function non-trivial.
(Bettini et al., 2024) use the Wasserstein metric for continuous policies (Vaserstein, 1969) as distance function $D$, while (McKee et al., 2022) use the total variation distance for discrete policies. For discrete policies, Wasserstein distance would require a cost function representing the cost of changing from one action to another, which might not be trivial to come up with. On the other hand, although well-suited for discrete policies, TVD might miss changes in action probabilities because it measures the largest difference assigned to an event (i.e. action) between two probability distributions.
We consider a simple example to illustrate this point. Suppose we have two policies $\pi^{1}$ and $\pi^{2}$ with action probabilities as shown in Figure 8. $\pi^{1}$ stays constant, while $\pi^{2}$ changes gradually over timesteps. We see that even as $\pi^{2}$ changes over time, the $TVD(\pi^{1},\pi^{2})$ between $\pi^{1}$ and $\pi^{2}$ remains constant. This is because TVD only measures the largest difference between the two distributions, and does not consider the overall difference between them. On the other hand, the Jensen-Shannon distance (JSD) (Endres & Schindelin, 2003), which is the square root of the Jensen-Shannon divergence, does not have this problem as it is a smooth distance function. Furthermore, it satisfies the conditions for being a metric – it is non-negative, symmetry, and it satisfies the triangle inequality.
For continuous policies, as shown in Figure 9, JSD exhibits similar trends to the Wasserstein distance. Since JSD is a reasonable metric for both continuous and discrete probability distributions, we will use it as the distance metric for all experiments and propose it as a suitable distance function for measuring policy diversity in MARL.
We also summarise the properties of the various distance metrics in Table 5.
Figure 8: Gradual changes in $\pi^{2}$, result in gradual changes in the Jensen-Shannon distance (JSD), while the Total Variation Distance (TVD) can miss changes in action probabilities.
Figure 9: Jensen-Shannon distance (JSD) trends similarly to Wasserstein distance when we have continuous policies.
Method Kinds of Actions Metric Smooth Formula
[…]
Total Variation Distance Discrete Metric No $TV(p,q)=\frac{1}{2}\sum_{x}|p(x)-q(x)|$
Jensen-Shannon Divergence (Lin, 1991) Both Divergence Yes $JSD(p\parallel q)=\frac{1}{2}D_{KL}(p\parallel m)+\frac{1}{2}D_{KL}(q\parallel m),\,m=\frac{1}{2}(p+q)$
Jensen-Shannon Distance (Endres & Schindelin, 2003) Both Metric Yes $\sqrt{JSD(p\parallel q)}$
Table 5: Measure Policy Diversity
Appendix B Specialisation and Synchronisation Game
Theorem B.1.
A shared policy cannot learn the optimal behaviour for the two-player Specialisation Game.
Proof.
Let $\pi$ be a shared policy for both agents, and let $\alpha=\mathbb{P}(a_{i}=0)$ represent the probability of any agent choosing action 0. The expected return of $\pi$ for each agent is:
$E[R(\pi)]$ $=\mathbb{P}(a_{0}=0,a_{1}=0)\cdot 0.5+\mathbb{P}(a_{0}=0,a_{1}=1)\cdot 1$ (10)
$\quad+\mathbb{P}(a_{0}=1,a_{1}=0)\cdot 1+\mathbb{P}(a_{0}=1,a_{1}=1)\cdot 0.5$ (11)
$=0.5\alpha^{2}+2\alpha(1-\alpha)+0.5(1-\alpha)^{2}$ (12)
$=-\alpha^{2}+\alpha+0.5$ (13)
$=-(\alpha-0.5)^{2}+0.75$ (14)
Thus, $E[R(\pi)]\leq 0.75<1$ for all $\alpha\in[0,1]$, with the maximum at $\alpha=0.5$. Therefore, a shared policy cannot achieve the optimal return of 1, confirming the need for specialised behaviour to optimise rewards. ∎
In Table 6, we present the results for the Specialisation and Synchronisation Games when actions are sampled during evaluation rather than selecting the argmax, as was done in Table 2.
Table 6: Average Evaluation Reward at the end of training (mean $\pm$ std) for REINFORCE variants in the Specialisation and Synchronisation Games, across 2, 4, and 8 Agents. Bold indicates values within 1 std of the maximum. Actions are sampled during evaluation.
Specialisation Game
Method 2 4 8
NoPS .9996 $\pm$ .0004 .9988 $\pm$ .0007 .9989 $\pm$ .0008
FuPS .7493 $\pm$ .0062 .6826 $\pm$ .0039 .6545 $\pm$ .0025
FuPS+ID .$9$$\pm$ .0002 .9995 $\pm$ .0003 .7275 $\pm$ .0293
Synchronisation Game
Method 2 4 8
NoPS .9993 $\pm$ .0004 .9976 $\pm$ .0022 .1250 $\pm$ 0
FuPS .$9$$\pm$ .0002 1.0 $\pm$ 0 .3000 $\pm$ .3500
FuPS+ID .$9$$\pm$ 0 .$9$$\pm$ .0003 .1250 $\pm$ 0
Appendix C HyperMARL Details
C.1 HyperMARL Pseudocode
In Algorithm 1, we present the pseudocode for HyperMARL, with HyperMARL-specific steps highlighted in blue. HyperMARL leverages hypernetworks to dynamically generate the parameters of both actor and critic networks. The weights of the hypernetworks and the agent embeddings are automatically updated through automatic differentiation (autograd) based on the computed loss. Additionally, Figure 1 provides a visual representation of the HyperMARL architecture.
Algorithm 1 HyperMARL
1:  Input: Number of agents $n$, number of training iterations $K$, MARL algorithm parameters (e.g., MAPPO-specific hyperparameters)
2:  Initialize:
3:   Hypernetwork parameters $\psi,\varphi$Ensure $\theta^{i}$ and $\phi^{i}$ follow standard initialization schemes, e.g., orthogonal}
4:   Agent embeddings $\{e^{i}\}_{i=1}^{n}$One-hot or orthogonally initialized learnable parameters}
5:  Output: Optimized joint policy $\boldsymbol{\pi}$
6:  for each training iteration $k=0,1,\dots,K-1$ do
7:     for each agent $i=1,\dots,n$ do
8:        $\theta^{i}\leftarrow h_{\psi}^{\pi}(e^{i})$Policy parameters}
9:        $\phi^{i}\leftarrow h_{\varphi}^{V}(e^{i})$Critic parameters}
10:     end for
11:     Interact with environment using $\{\pi_{{\color[rgb]{0.2734375,0.51171875,0.70703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.51171875,0.70703125}\theta^{i}}}\}_{i=1}^{n}$ to collect trajectories $\mathcal{D}$
12:     Compute shared loss $\mathcal{L}$ from $\mathcal{D}$, using $\{V_{{\color[rgb]{0.2734375,0.51171875,0.70703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.51171875,0.70703125}\phi^{i}}}\}_{i=1}^{n}$Standard RL loss function}
13:     Update $\psi$, $\varphi$, and $\boldsymbol{e}$ by minimizing $\mathcal{L}$Optimise parameters.}
14:  end for
15:  Return $\boldsymbol{\pi}=(\pi^{1},\dots,\pi^{n})$
C.2 Scalability and Parameter Efficiency
Hypernetworks generate weights for the target network, which can lead to high-dimensional outputs and many parameters for deep target networks. This challenge is amplified in MLP-based hypernetworks, which include additional hidden layers. Figure 10 shows scaling trends:
  • NoPS and linear hypernetworks: Parameter count grows linearly with the number of agents.
  • FuPS: More efficient, as growth depends on one-hot vector size.
  • MLP hypernetworks: Scale better with larger populations, since they only require embeddings of fixed size for each new agent.
To reduce parameter count, techniques like shared hypernetworks, chunked hypernetworks (von Oswald et al., 2020; Chauhan et al., 2024), or producing low-rank weight approximations, can be used. While naive implementations are parameter-intensive, this might be less critical in RL and MARL which commonly have smaller actor-critic networks. Moreover, HyperMARL’s near-constant scaling with agents suggests strong potential for large-scale MARL applications.
Figure 10: Parameter scaling for IPPO variants with increasing agents (4 to 1024). MLP Hypernetworks scale nearly constantly, while NoPS, Linear Hypernetworks, and FuPS+One-Hot grow linearly. Log scale on both axes.
Figure 11: Dispersion performance with four agents. FuPS variants match HyperMARL in parameter count but still underperform.
To isolate the effects of parameter count, we scaled the FuPS networks (Figure 11) to match the number of trainable parameters in HyperMARL. Despite generating 10x smaller networks, HyperMARL consistently outperforms FuPS variants, showing its advantages extend beyond parameter count.
Appendix D Experiment Details
D.1 Environments
(a) Dispersion.
(b) Navigation.
(c) MAMuJoCo.
(d) SMAX.
Figure 12: Multi-Agent environments used in our experiments.
Dispersion (VMAS) (Bettini et al., 2022): A 2D environment where four agents collect unique food particles. This task requires specialised heterogeneous behaviours and resembles the Specialisation Game from Section 3.1.
Navigation (VMAS) (Bettini et al., 2022): Agents navigate in a 2D space to assigned goals, receiving dense rewards based on proximity. Configurations include shared goals (homogeneous), unique goals (heterogeneous), and mixed goals, where some agents share goals while others have unique ones.
Multi-Agent MuJoCo (MAMuJoCo) (Peng et al., 2021): A multi-agent extension of MuJoCo, where robot body parts (e.g., a cheetah’s legs) are modelled as different agents. Agents coordinate to perform efficient motion, requiring heterogeneous policies (Zhong et al., 2024).
SMAX (JaxMARL) (Rutherford et al., 2024): Discrete action tasks with 2 to 20 agents on SMACv1- and SMACv2-style maps. FuPS baselines have been shown optimal for these settings (Yu et al., 2022; Fu et al., 2022) indicating homogeneous behaviour is preferred here.
Table 7: Overview of Baseline Methods and Justifications. Selected baselines (✓) are chosen based on their relevance to our focus on parameter sharing and specialisation/generalisation in MARL. Non-selected methods (✗) are excluded due to implementation limitations or misalignment with our scope.
Baseline Method Category Chosen Justification
IPPO (NoPS, FuPS) (De Witt et al., 2020) NoPS/FuPS Serves as a strong MARL baseline for both independent (NoPS) and shared (FuPS) policies.
MAPPO (NoPS, FuPS) (Yu et al., 2022) NoPS/FuPS Provides a robust MARL baseline with centralized critics for both NoPS and FuPS configurations.
DiCo (Bettini et al., 2024) Architectural Enables comparison with a method that employs preset diversity levels and maintains both non-shared and shared parameters.
HAPPO (Zhong et al., 2024) Sequential Updates Facilitates comparison with a method designed for heterogeneous behaviours, utilizing sequential updates and maintaining non-shared parameters per agent.
SEAC (Christianos et al., 2020) Shared Experience Focuses on experience sharing rather than parameter sharing, which is outside our scope.
SePS (Christianos et al., 2021) SePS Requires pretraining, which is beyond the scope of our current study.
CDAS (Li et al., 2021) Intrinsic Reward Only implemented for off-policy methods which are outside our scope, and has been shown to underperform FuPS/NoPS (Fu et al., 2022).
ROMA/RODE (Wang et al., 2020b, a) Role-based Demonstrates limited practical performance (Christianos et al., 2020).
SNP-PS (Kim & Sung, 2023) Architectural/Pruning No available code implementation.
Kaleidoscope (Li et al., 2024) Architectural/Pruning Only implemented for off-policy methods, which is outside our scope.
D.1.1 Training and Evaluation
  • Training:
    • For Dispersion (5.2), we run 5 seeds and train for 20 million timesteps.
    • For Navigation (5.2), SMAX (5.3), and MaMuJoCo (5.2), we run 5 seeds and train for 10 million timesteps, consistent with the baselines.
  • Evaluation:
    • For Dispersion (5.2), evaluation is performed every 100k timesteps across 32 episodes.
    • For Navigation (5.2), following the baselines, evaluation is performed every 120k timesteps across 200 episodes.
    • For SMAX (5.3), evaluation is performed every 500k timesteps across 32 episodes.
    • For MaMuJoCo (5.2), following the baselines, evaluation is performed every 25 training episodes over 40 episodes.
D.1.2 Measuring Policy Diversity Details
We measure team diversity using the System Neural Diversity (SND) metric (Equation 9 (Bettini et al., 2023), details Section A) with Jensen-Shannon distance. SND ranges from 0 (identical policies across all agents) to 1 (maximum diversity). We collect a dataset of observations from IPPO-NoPS and IPPO-FuPS policies checkpointed at 5 and 20 million training steps. Each policy is rolled out for 10,000 episodes, generating 16 million observations. We then sample 1 million observations from this dataset to calculate the SND for each method tested.
D.2 HyperMARL Architecture
For the experiments in Sections 5.2 and 5.2, we use feedforward architectures, while for Section 5.3, we use recurrent GRU networks (Cho, 2014). For recurrent PPO variants, HyperMARL maintains the same architecture as for feedforward methods, generating only the actor and critic feedforward weights, not the GRU weights.
Appendix E Detailed Results
E.1 Dispersion Detailed Results
E.1.1 Sample Efficiency Plots
(a) IPPO Sample Efficiency
(b) MAPPO Sample Efficiency
Figure 13: Sample efficiency of IPPO and MAPPO variants on the Dispersion environment. While FuPS methods struggle to converge to optimal performance, Hypernetworks achieve efficiency comparable to NoPS. Even when running MAPPO-FuPS for longer (see Figure 14), it converges to suboptimal performance.
Figure 14: We see that even if we run MAPPO-FuPS on Dispersion for 40 million timesteps (double the timesteps of MLP Hypernetwork), it converges to suboptimal performance.
E.2 Detailed MAMujoco Plots
(a) Humanoid-v2 17x1
(b) Walker2d-v2 2x3
(c) HalfCheetah-v2 2x3
(d) Ant-v2 4x2
Figure 15: Performance of Recurrent IPPO and MAPPO on SMAX. HyperMARL performs comparably to these baselines.
E.3 Detailed Navigation Plots
(a) Same Goal (8 agents)
(b) Different Goals (8 agents)
(c) Four Goals (8 agents)
(d) Scenarios with 8 Agents
(e) Same Goal (4 agents)
(f) Different Goals (4 agents)
(g) Two Goals (4 agents)
(h) Scenarios with 4 Agents
(i) Same Goal (2 agents)
(j) Different Goals (2 agents)
(k) Scenarios with 2 Agents
Figure 17: Sample Efficiency of IPPO variants in the VMAS Navigation environment under different goal configurations and agent numbers.
E.4 Additional Ablations
Figure 18: Ablation results comparing HyperMARL with its variants in Dispersion. The results highlight that gradient decoupling is essential for maintaining HyperMARL’s performance.
Appendix F Hyperparameters
Table 8: Hyperparameters for Specialisation and Synchronisation Game
Hyperparameter Value
Number of agents and foods 2, 4, 8
Number of foods 3
Number of seeds 5
Number of trials 10,000
Evaluation trials 1,000
Batch size 10
Learning rate 0.01
Activation function ReLU
Output activation Softmax
Optimizer SGD
Total Number of Parameters
   2 Agents
NoPS: 36
FuPS: 34
FuPS+ID: 42
   4 Agents
NoPS: 112
FuPS: 100
FuPS+ID: 148
   8 Agents
NoPS: 384
FuPS: 328
FuPS+ID: 552
Table 9: IPPO and MAPPO Hyperparameters in Dispersion
Hyperparameter Value
LR 0.0005
GAMMA 0.99
VF_COEF 0.5
CLIP_EPS 0.2
ENT_COEF 0.01
NUM_ENVS 16
NUM_STEPS 128
GAE_LAMBDA 0.95
NUM_UPDATES 9765
EVAL_EPISODES 32
EVAL_INTERVAL 100000
MAX_GRAD_NORM 0.5
UPDATE_EPOCHS 4
NUM_MINIBATCHES 2
TOTAL_TIMESTEPS 20000000
ANNEAL_LR false
ACTOR_LAYERS [64, 64]
CRITIC_LAYERS [64, 64]
ACTIVATION relu
SEEDS 30,1,42,72858,2300658
ACTION_SPACE_TYPE discrete
Table 10: MLP Hypernet Hyperparameters in Dispersion
Parameter IPPO MAPPO
HYPERNET_EMBEDDING_DIM 4 8
EMBEDDING_DIM Sweep [4, 16, 64] [4, 8, 16, 64]
HYPERNET_HIDDEN_DIMS 64 64
Table 11: Dispersion Settings
Setting Value
n_food 4
n_agents 4
max_steps 100
food_radius 0.08
share_reward false
penalise_by_time true
continuous_actions false
Table 12: IPPO Hyperparameters for Navigation
Hyperparameters Value
LR 0.00005
NUM_ENVS 600
NUM_STEPS 100
TOTAL_TIMESTEPS $10^{6}$
UPDATE_EPOCHS 45
NUM_MINIBATCHES 30
GAMMA 0.9
GAE_LAMBDA 0.9
CLIP_EPS 0.2
ENT_COEF 0.0
VF_COEF 1.0
MAX_GRAD_NORM 5
ACTIVATION tanh
ANNEAL_LR False
ACTOR_LAYERS [256, 256]
CRITIC_LAYERS [256, 256]
ACTION_SPACE_TYPE continuous
Table 13: MLP Hypernet Hyperparameters in Navigation
Parameter IPPO MAPPO
HYPERNET_EMBEDDING_DIM 4 8
EMBEDDING_DIM Sweep [4, 16, 64] [4, 8, 16, 64]
HYPERNET_HIDDEN_DIMS 64 64
Table 14: DiCo Algorithm $SND\_des$ Hyperparameter
Goal Configuration Number of Agents SND_des
All agents same goal 2 0
4 0
8 0
All agents different goals 2 1.2 (From DiCo paper)
4 [-1,1.2,2.4] $\Rightarrow$ -1 (Best)
8 [-1,1.2,4.8] $\Rightarrow$ -1 (Best)
Some agents share goals 4 [-1,1.2] $\Rightarrow$ -1 (Best)
8 [-1,2.4,1.2] $\Rightarrow$ -1 (Best)
Table 15: Parameter Sweeps for IPPO Variants in Navigation Tasks with Four and Eight Agents
Parameter Sweeps
CLIP_EPS 0.2, 0.1
LR 5e-5, 5e-4, 2.5e-4
Algorithm Setting Selected Values
IPPO-FuPS 8 Agents (Same Goals) 0.2, 5e-5
8 Agents (Different Goals) 0.1, 5e-5
8 Agents (Four Goals) 0.1, 5e-5
4 Agents (Same Goals) 0.2, 5e-5
4 Agents (Different Goals) 0.2, 5e-5
4 Agents (Two Goals) 0.2, 5e-5
IPPO-Linear Hypernetwork 8 Agents (Same Goals) 0.2, 5e-5
8 Agents (Different Goals) 0.1, 5e-5
8 Agents (Four Goals) 0.1, 5e-5
4 Agents (Same Goals) 0.2, 5e-5
4 Agents (Different Goals) 0.1, 5e-5
4 Agents (Two Goals) 0.1, 5e-5
IPPO-MLP Hypernetwork 8 Agents (Same Goals) 0.2, 5e-5
8 Agents (Different Goals) 0.1, 5e-5
8 Agents (Four Goals) 0.1, 5e-5
4 Agents (Same Goals) 0.1, 5e-5
4 Agents (Different Goals) 0.1, 5e-5
4 Agents (Two Goals) 0.1, 5e-5
IPPO-NoPS 8 Agents (Same Goals) 0.1, 5e-5
8 Agents (Different Goals) 0.2, 5e-5
8 Agents (Four Goals) 0.1, 5e-5
4 Agents (Same Goals) 0.1, 5e-5
4 Agents (Different Goals) 0.2, 5e-5
4 Agents (Two Goals) 0.1, 5e-5
IPPO-Dico 8 Agents (Same Goals) 0.2, 5e-5
8 Agents (Different Goals) 0.1, 2.5e-4
8 Agents (Four Goals) 0.1, 2.5e-4
4 Agents (Same Goals) 0.2, 5e-5
4 Agents (Different Goals) 0.1, 2.5e-4
4 Agents (Two Goals) 0.1, 5e-4
Table 16: Environment Settings for Navigation Task
Parameter Value
n_agents 2,4,8
agents_with_same_goal 1, n_agents/2, n_agents
max_steps 100
collisions False
split_goals False
observe_all_goals True
shared_rew False
lidar_range 0.35
agent_radius 0.1
continuous_actions True
Table 17: Default algorithm and model hyperparameters for the Ant-v2-4x2 environment (from  (Zhong et al., 2024)).
Parameter Value
— Algorithm Parameters —
action_aggregation prod
actor_num_mini_batch 1
clip_param 0.1
critic_epoch 5
critic_num_mini_batch 1
entropy_coef 0
fixed_order true
gae_lambda 0.95
gamma 0.99
huber_delta 10.0
max_grad_norm 10.0
ppo_epoch 5
share_param false
use_clipped_value_loss true
use_gae true
use_huber_loss true
use_max_grad_norm true
use_policy_active_masks true
value_loss_coef 1
— Model Parameters —
activation_func relu
critic_lr 0.0005
data_chunk_length 10
gain 0.01
hidden_sizes [128, 128, 128]
initialization_method orthogonal_
lr 0.0005
opti_eps 1e-05
recurrent_n 1
std_x_coef 1
std_y_coef 0.5
use_feature_normalization true
use_naive_recurrent_policy false
use_recurrent_policy false
weight_decay 0
Table 18: Default algorithm and model hyperparameters for the Humanoid-v2-17x1 environment (from  (Zhong et al., 2024)).
Parameter Value
— Algorithm Parameters —
action_aggregation prod
actor_num_mini_batch 1
clip_param 0.1
critic_epoch 5
critic_num_mini_batch 1
entropy_coef 0
fixed_order false
gae_lambda 0.95
gamma 0.99
huber_delta 10.0
max_grad_norm 10.0
ppo_epoch 5
share_param false
use_clipped_value_loss true
use_gae true
use_huber_loss true
use_max_grad_norm true
use_policy_active_masks true
value_loss_coef 1
— Model Parameters —
activation_func relu
critic_lr 0.0005
data_chunk_length 10
gain 0.01
hidden_sizes [128, 128, 128]
initialization_method orthogonal_
lr 0.0005
opti_eps 1e-05
recurrent_n 1
std_x_coef 1
std_y_coef 0.5
use_feature_normalization true
use_naive_recurrent_policy false
use_recurrent_policy false
weight_decay 0
Table 19: Default algorithm and model hyperparameters for the Walker2d-v2-2x3 environment (from  (Zhong et al., 2024)).
Parameter Value
— Algorithm Parameters —
action_aggregation prod
actor_num_mini_batch 2
clip_param 0.05
critic_epoch 5
critic_num_mini_batch 2
entropy_coef 0
fixed_order false
gae_lambda 0.95
gamma 0.99
huber_delta 10.0
max_grad_norm 10.0
ppo_epoch 5
share_param false
use_clipped_value_loss true
use_gae true
use_huber_loss true
use_max_grad_norm true
use_policy_active_masks true
value_loss_coef 1
— Model Parameters —
activation_func relu
critic_lr 0.001
data_chunk_length 10
gain 0.01
hidden_sizes 128, 128, 128
initialization_method orthogonal_
lr 0.001
opti_eps 1e-05
recurrent_n 1
std_x_coef 1
std_y_coef 0.5
use_feature_normalization true
use_naive_recurrent_policy false
use_recurrent_policy false
weight_decay 0
Table 20: Default algorithm and model hyperparameters for the HalfCheetah-v2-2x3 environment (from  (Zhong et al., 2024)).
Parameter Value
— Algorithm Parameters —
action_aggregation prod
actor_num_mini_batch 1
clip_param 0.05
critic_epoch 15
critic_num_mini_batch 1
entropy_coef 0.01
fixed_order false
gae_lambda 0.95
gamma 0.99
huber_delta 10.0
max_grad_norm 10.0
ppo_epoch 15
share_param false
use_clipped_value_loss true
use_gae true
use_huber_loss true
use_max_grad_norm true
use_policy_active_masks true
value_loss_coef 1
— Model Parameters —
activation_func relu
critic_lr 0.0005
data_chunk_length 10
gain 0.01
hidden_sizes 128, 128, 128
initialization_method orthogonal_
lr 0.0005
opti_eps 1e-05
recurrent_n 1
std_x_coef 1
std_y_coef 0.5
use_feature_normalization true
use_naive_recurrent_policy false
use_recurrent_policy false
weight_decay 0
Table 21: HyperMARL Hyperparameters Across MaMuJoCo Environments
Parameter Humanoid-v2-17x1 Walker2d-v2-2x3 HalfCheetah-v2-2x3 Ant-v2-4x2 Sweeps
AGENT_ID_EMBEDDING_DIM 64 64 64 64 None
HYPERNET_HIDDEN_DIMS 64 64 64 64 None
clip_param 0.075 0.0375 0.0375 0.075 [0.1,0.075,0.05,0.0375]
Table 22: Recurrent IPPO and MAPPO Hyperparameters in SMAX (from JaxMARL paper)
Hyperparameter IPPO Value MAPPO Value
LR 0.004 0.002
NUM_ENVS 128 128
NUM_STEPS 128 128
GRU_HIDDEN_DIM 128 128
FC_DIM_SIZE 128 128
TOTAL_TIMESTEPS 1e7 1e7
UPDATE_EPOCHS 4 4
NUM_MINIBATCHES 4 4
GAMMA 0.99 0.99
GAE_LAMBDA 0.95 0.95
CLIP_EPS 0.05 0.2
SCALE_CLIP_EPS False False
ENT_COEF 0.01 0.0
VF_COEF 0.5 0.5
MAX_GRAD_NORM 0.25 0.25
ACTIVATION relu relu
SEED 30,1,42,72858,2300658 30,1,42,72858,2300658
ANNEAL_LR True True
OBS_WITH_AGENT_ID - True
Table 23: Hyperparameter Sweeps and Final Values for Different Maps in SMAX. H- refers to HyperMARL MLP Hypernetworks.
Map Algorithm LR Range Chosen LR HNET Embedding Dim HNET Hidden Dims
2s3z IPPO 0.004 0.004 -
MAPPO 0.002 0.002 - -
H-IPPO 0.004 0.004 4 32
H-MAPPO 0.002 0.002 64 16
3s5z IPPO 0.004 0.004 - -
MAPPO 0.002,0.005,0.0003 0.002 - -
H-IPPO 0.004 0.004 64 16
H-MAPPO 0.002,0.005,0.0003 0.0003 64 16
smacv2_10_units IPPO 0.005,0.001,0.0003,0.004 0.001 - -
MAPPO 0.002,0.005,0.0003 0.0003 - -
H-IPPO 0.005,0.001,0.0003,0.004 0.005 4 64
H-MAPPO 0.002,0.005,0.0003 0.0003 64 16
smacv2_20_units IPPO 0.002,0.005,0.0003 0.005 - -
MAPPO 0.002,0.005,0.0003 0.0003 - -
H-IPPO 0.002,0.005,0.0003 0.005 64 64
H-MAPPO 0.002,0.005,0.0003 0.0003 4 64
Note: HNET Embedding Dim refers to the hypernetwork embedding dimension, chosen from the range {4, 16, 64}. HNET Hidden Dims refers to the hidden layer dimensions of the hypernetwork, chosen from the range {16, 32, 64}.