### v3

# HyperMARL: Adaptive Hypernetworks for Multi-Agent RL

Abstract
Adaptability to specialised or homogeneous behaviours is critical in cooperative multi-agent reinforcement learning (MARL). Parameter sharing (PS) techniques, common for efficient adaptation, often limit behavioural diversity due to cross-agent gradient interference, which we show can be exacerbated by the coupling of observations and agent IDs. Current remedies typically add complexity through altered objectives, manual preset diversity levels, or sequential updates. We ask: can shared policies adapt without these complexities? We propose HyperMARL, a PS approach using hypernetworks for dynamic agent-specific parameters, without altering the RL objective or requiring preset diversity levels. HyperMARL’s explicit decoupling of observation- and agent-conditioned gradients empirically reduces policy gradient variance, facilitates shared-policy adaptation (including specialisation), and helps mitigate cross-agent interference. Across diverse MARL benchmarks (up to 20 agents), requiring homogeneous, heterogeneous, or mixed behaviours, HyperMARL achieves competitive performance against key baselines – fully shared, non-parameter sharing, and three diversity-promoting methods – while preserving behavioural diversity comparable to non-parameter sharing. These findings establish HyperMARL as a versatile approach for adaptive MARL. The code is publicly available at https://github.com/KaleabTessera/HyperMARL.
1 Introduction
Specialist and generalist behaviours are critical to collective intelligence, enhancing performance and adaptability in both natural and artificial systems [47, 38, 39, 22, 45]. In Multi-Agent Reinforcement Learning (MARL), this translates to a critical need for policies that can flexibly exhibit specialised, homogeneous, or mixed behaviours to meet diverse task demands [25, 6, 2].
Figure 1: HyperMARL Policy Architecture. Common Agent-ID conditioned shared MARL policy (left) vs HyperMARL (right), which uses a hypernetwork to generate agent-specific policies and decouples observation- and agent-conditioned gradients.
Optimal MARL performance thus hinges on being able to represent the required behaviours. While No Parameter Sharing (NoPS) [28] enables specialisation by using distinct per-agent networks, it suffers from significant computational overhead and sample inefficiency [10]. Conversely, Full Parameter Sharing (FuPS) [40, 17, 14], which trains a single shared network, improves efficiency but typically struggles to foster the behavioural diversity necessary for many complex tasks [23, 16, 25].
Balancing FuPS efficiency with the capacity for diverse behaviours therefore remains a central open problem in MARL. Prior works have explored intrinsic-rewards [25, 21], role-based allocations [43, 44], specialised architectures [23, 26, 6], or sequential updates [49]. However, these methods often alter the learning objective, require prior knowledge of optimal diversity levels, involve delicate hyperparameter tuning, or necessitate maintaining agent-specific parameters and sequential updates (see Section 7).
These challenges lead us to ask: Can we design a shared MARL architecture that flexibly supports both specialised and homogeneous behaviours—without altered learning objectives, manual preset diversity levels, or sequential updates? A key difficulty with FuPS, particularly for diverse behaviours, was hypothesised to be gradient interference among agents, whereby their updates negatively impact each other’s learning [10, 49]. We not only empirically validate this hypothesis but also demonstrate a critical insight: this conflict is significantly exacerbated by the common practice of coupling observations with agent IDs within a shared network (Fig. 1 for coupling, Sec. 3.2 for results).
To counteract the challenges posed by this coupling, we propose HyperMARL, a novel parameter sharing method that explicitly decouples observation- and agent-conditioned gradients (Section 4.3) through hypernetworks [18]. Specifically, agent-conditioned hypernetworks, which condition on an agent identifier (or embedding), and generate an agent’s actor and critic weights on the fly (Fig. 1). Hypernetworks have already proved effective for resolving gradient conflicts in multi-task RL [30] and for mitigating catastrophic forgetting in continual learning [42], our work establishes their effectiveness in MARL. We further show that HyperMARL empirically attains lower policy gradient variance than FuPS and that this decoupling is critical for specialisation (Sec. 5.26), suggesting it helps mitigate cross-agent gradient interference in shared architectures.
We validate HyperMARL on diverse MARL benchmarks—including Dispersion and Navigation (VMAS) [5], Multi-Agent MuJoCo (MAMuJoCo) [34], and SMAX [36]—across environments with two to twenty agents that require homogeneous, heterogeneous, or mixed behaviours. HyperMARL consistently matches or outperforms NoPS, FuPS, and diversity-promoting methods such as DiCo [6],HAPPO [49] and Kaleidoscope [26], while achieving NoPS-level behavioural diversity while using a shared architecture.
Our contributions are as follows:
  • Motivated by evidence that coupling agent IDs with observations can be detrimental, we propose HyperMARL (Sec. 4), a method that leverages agent-conditioned hypernetworks to decouple observation- and agent-conditioned gradients. HyperMARL adaptively learns diverse, homogeneous, or mixed behaviours without altering the RL objective or requiring preset diversity levels.
  • We thoroughly evaluate HyperMARL (Section 5) across diverse MARL environments with up to 20 agents. HyperMARL consistently achieves competitive mean episode returns compared to NoPS, FuPS, and diversity-focused methods such as DiCo, HAPPO and Kaleidoscope, while achieving NoPS-level behavioural diversity.
  • We demonstrate that HyperMARL empirically reduces policy gradient variance compared to FuPS and facilitates specialisation, suggesting the importance of gradient decoupling for mitigating cross-agent interference (Section 5.2; Section 6).
2 Background
We formulate the fully cooperative multi-agent systems addressed in our work as a Dec-POMDP [31]. A Dec-POMDP is a tuple, $\langle\mathbb{I},\mathbb{S},\{\mathbb{A}^{i}\}_{i\in\mathbb{I}},R,\{\mathbb{O}^{i}\}_{i\in\mathbb{I}},O,T,\rho_{0},\gamma\rangle$, where $\mathbb{I}$ is the set of agents of size $n=|\mathbb{I}|$, $\mathbb{S}$ is the set of global states with an initial state distribution $\rho_{0}$, $\mathbb{A}^{i}$ is the action space for agent $i$ where $\mathbb{A}=\times_{i}\mathbb{A}^{i}$ is the joint action space, $R:\mathbb{S}\times\mathbb{A}\to\mathbb{R}$ is the shared reward function, $\mathbb{O}^{i}$ is the observation space for agent $i$ with the joint observation space $\mathbb{O}=\times_{i}\mathbb{O}^{i}$, $O:\mathbb{O}\times\mathbb{A}\times\mathbb{S}\to[0,1]$ is the probability of joint observation $\mathbf{o}\in\mathbb{O}$, i.e. $O(\mathbf{o},\mathbf{a},s)=\Pr(\mathbf{o_{t}}|s_{t},\mathbf{a_{t-1}})$, $T:\mathbb{S}\times\mathbb{A}\times\mathbb{S}\to[0,1]$ is the state transition function i.e. $T\left(s,\mathbf{a},s^{\prime}\right)=\Pr\left(s_{t+1}\mid s_{t},\mathbf{a}_{t}\right)$ and $\gamma$ is the discount factor.
In this setting, each agent $i$ receives a partial observation $o_{t}^{i}\in\mathbb{O}^{i}$. These observations are accumulated into an action-observation history $h_{t}^{i}=\left(o_{0}^{i},a_{0}^{i},\ldots,o_{t-1}^{i},a_{t-1}^{i},o_{t}^{i}\right)$. Each agent $i$ acts based on their decentralized policies $\pi^{i}(a^{i}|h^{i})$. The joint history and joint action are defined as follows $\mathbf{h_{t}}=(h_{t}^{1},\ldots,h_{t}^{n})$ and $\mathbf{a_{t}}=(a_{t}^{1},\ldots,a_{t}^{n})$. The goal is to learn an optimal joint policy $\boldsymbol{\pi^{*}}=(\pi^{1*},...,\pi^{n*})$ that maximizes the expected discounted return as follows, [^1] We use simplified notion here omitting explicit dependence on state transitions and distributions for brevity. $\boldsymbol{\pi}^{*}=\arg\max_{\boldsymbol{\pi}}\mathbb{E}_{s_{0}\sim\rho_{0},\;\mathbf{h}\sim\boldsymbol{\pi}}\left[G(\mathbf{h})\right]$, where $G(\mathbf{h})=\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},\mathbf{a}_{t})$.
Specialised Policies and Environments. We say an environment is specialised if its optimal joint policy contains at least two distinct, non-interchangeable agent policies ( Def. 1 in App. C). Under this mild condition, tasks such as Dispersion (5.2) or our Specialisation Game (E.1) require agents to learn complementary roles rather than identical behaviours.
3 Are Independent or Fully Shared Policies Enough?
Standard independent (NoPS) and fully parameter-shared (FuPS) policies face inherent trade-offs in MARL. NoPS allows for uninhibited agent specialisation but can be sample inefficient and computationally expensive. FuPS, often conditioned with an agent ID (FuPS+ID), is more efficient but can struggle when agents must learn diverse behaviours. This section investigates the limitations of these common policy architectures.
P2
A B
P1 A (0.5,0.5) (1,1)
B (1,1) (0.5,0.5)
Two-Player Payoff matrix
$N$-player Interaction
(a) Specialisation Game
P2
A B
P1 A (1,1) (0.5,0.5)
B (0.5,0.5) (1,1)
Two-Player Payoff matrix
$N$-player Interaction
(b) Synchronisation Game
Figure 2: Specialisation and Synchronisation Games. The Specialisation game (left), which encourages distinct actions, and the Synchronisation game (right), where rewards encourage identical actions. Depicted are their two-player payoff matrices (pure Nash equilibria in blue) and $N$-player interaction schematics. While simple in form, these games are challenging MARL benchmarks due to non-stationarity and exponentially scaling observation spaces (temporal version).
To probe these limitations, we introduce two illustrative environments: the Specialisation Game, rewarding distinct actions, and the Synchronisation Game, rewarding identical actions. Both are inspired by prior work [16, 5, 32] and extended here to $N$-agent and temporal settings where agents observe prior joint actions (see Appendix E for full definitions).
3.1 Limitations of Fully Shared and Independent Policies
Figure 3: Average evaluation reward (mean $\pm$ 95% CI) for temporal Specialisation vs. Synchronisation using REINFORCE (10 seeds). Bold: highest mean, no CI overlap. Neither fully shared nor independent policies consistently achieve the highest mean reward.
Specialisation Synchronisation
#Ag NoPS FuPS FuPS+ID NoPS FuPS FuPS+ID
2 0.88$\pm$0.09 0.50$\pm$0.00 0.64$\pm$0.10 0.83$\pm$0.12 1.00$\pm$0.00 0.91$\pm$0.09
4 0.74$\pm$0.08 0.25$\pm$0.00 0.40$\pm$0.07 0.32$\pm$0.03 1.00$\pm$0.00 0.67$\pm$0.15
8 0.68$\pm$0.02 0.12$\pm$0.00 0.25$\pm$0.03 0.14$\pm$0.00 1.00$\pm$0.00 0.54$\pm$0.10
16 0.64$\pm$0.01 0.06$\pm$0.00 0.13$\pm$0.02 0.07$\pm$0.00 1.00$\pm$0.00 0.55$\pm$0.14
(a) Avg. evaluation reward
(b) Avg. gradient conflict
Figure 4: Multi-agent policy gradient methods in the Specialisation environment. The FuPS+ID (No State) ablation outperforms FuPS+ID, showing near-orthogonal gradients (4(b)), indicating that observation–ID decoupling is important. HyperMARL (MLP) enables this decoupling while leveraging state information, and achieves better performance and reduced gradient conflict than FuPS+ID.
FuPS without agent IDs provably cannot recover optimal pure Nash equilibria in the non-temporal 2-player Specialisation Game (Proof E.3, App. E). In practice, however, FuPS is often conditioned with agent IDs, and MARL policies must handle complexities beyond static, two-player interactions. We therefore evaluate standard architectures in the temporal $n$-player versions of these games [^2] Results for non-temporal (normal-form) variants are in App. E.5. We compare three standard architectures trained with REINFORCE [46]: 1) NoPS: independent policies ($\pi_{\theta^{i}}(a^{i}|o^{i})$); 2) FuPS: a single shared policy ($\pi_{\theta}(a^{i}|o^{i})$); and 3) FuPS+ID: a shared policy incorporating a one-hot agent ID ($\pi_{\theta}(a^{i}|o^{i},id^{i})$). All use single-layer networks, 10-step episodes, and $10,000$ training steps (further details in App. 6).
Empirical Performance. Table 3 shows that neither NoPS nor FuPS consistently achieves the highest mean evaluation rewards. NoPS excels in the Specialisation Game but is outperformed by FuPS (optimal) and FuPS+ID in the Synchronisation Game. Furthermore, the performance gaps widen as the number of agents increases (notably at $n=8$ and $n=16$), highlighting the scalability challenges of both fully independent and fully shared policies.
3.2 Why FuPS+ID Fails to Specialise: The Problem of Gradient Conflict
Despite being a universal approximator [19], FuPS+ID often struggles to learn diverse policies in practice (Table 3, [10, 49]). A key reason is gradient conflict: when a single network processes both observation $o$ and agent ID $id^{i}$, updates intended to specialise agent $i$ (based on $id^{i}$) can conflict with updates for agent $j$ (based on $id^{j}$), particularly if they share similar observations but require different actions. This obstructs the emergence of specialised behaviours (conflict measured via inter-agent gradient cosine similarity, App. E.4).
Importance of Observation and ID Decoupling To investigate the effect of entangled observation and ID inputs, we introduce an ablation: FuPS+ID (No State), where the policy $\pi_{\theta}(a^{i}\mid id^{i})$ conditions only on the agent ID, ignoring observations. Surprisingly, FuPS+ID (No State) outperforms standard FuPS+ID in the Specialisation Game for all tested $N$ (Figure 4(a)), even when $N\leq 4$ (where observation spaces are small, suggesting the issue is not merely observation size). Figure 4(b) reveals why: FuPS+ID (No State) shows near-zero gradient conflict (nearly orthogonal gradients), whereas standard FuPS+ID exhibits negative cosine similarities (conflicting gradients).
These results show that naively coupling observation and ID inputs in shared networks can lead to destructive interference, hindering specialisation. While discarding observations is not a general solution (most tasks require state information), this finding motivates designing architectures that can leverage both state and agent IDs, while minimising interference. Section 4 introduces HyperMARL (Figure 1), which explicitly decouples observation- and agent-conditioned gradients through agent-conditioned hypernetworks, leading to improved performance over FuPS variants and reduced gradient conflict compared to standard FuPS+ID (Figure 4).
4 HyperMARL
We introduce HyperMARL, an approach that uses agent-conditioned hypernetworks to learn diverse or homogeneous policies end-to-end, without modifying the standard RL objective or requiring manual preset diversity levels. By operating within a fully shared paradigm, HyperMARL leverages shared gradient information while enabling specialisation through the decoupling of observation- and agent-conditioned gradients. We present the pseudocode in Sec. F.1, with additional scaling (F.3) and runtime (F.4) details.
4.1 Hypernetworks for MARL
As illustrated in Figure 1, for any agent $i$ with context $e^{i}$ (i.e., either a one-hot encoded ID or a learned embedding), the hypernetworks generate the agent-specific parameters:
$\theta^{i}=h_{\psi}^{\pi}(e^{i}),\quad\phi^{i}=h_{\varphi}^{V}(e^{i}),$ (1)
where $h_{\psi}^{\pi}$ and $h_{\varphi}^{V}$ are the hypernetworks for the policy and critic, respectively. The parameters $\theta^{i}$ and $\phi^{i}$ define each agent’s policy $\pi_{\theta^{i}}$ and critic $V_{\phi^{i}}$, dynamically enabling either specialised or homogeneous behaviours as required by the task.
Linear Hypernetworks Given a one-hot agent ID, $\mathds{1}^{i}\in\mathbb{R}^{1\times n}$, a linear hypernetwork $h_{\psi}^{\pi}$ generates agent-specific parameters $\theta^{i}$ as :
$\theta^{i}=h_{\psi}^{\pi}(\mathds{1}^{i})=\mathds{1}^{i}\cdot W+b$ (2)
where $W\in\mathbb{R}^{n\times m}$ is the weight matrix (with $m$ the per-agent parameter dimensionality and $n$ is the number of agents), and $b\in\mathbb{R}^{1\times m}$ is the bias vector. Since $\mathds{1}^{i}$ is one-hot encoded, each $\theta^{i}$ corresponds to a specific row of $W$ plus the shared bias $b$. If there is no shared bias term, this effectively replicates training of separate policies for each task (in our case, for each agent) [3], since there are no shared parameters and gradient updates are independent.
MLP Hypernetworks for Expressiveness To enhance expressiveness, MLP Hypernetworks incorporate hidden layers and non-linearities:
$\theta^{i}=h_{\psi}^{\pi}(e^{i})=f_{\psi_{1}}^{\pi}\bigl{(}g_{\psi_{2}}^{\pi}(e^{i})\bigr{)}$ (3)
where $g_{\psi_{2}}^{\pi}$ is an MLP processing the agent context $e^{i}$, and $f_{\psi_{1}}^{\pi}$ is a final linear output layer.
Unlike linear hypernetworks with one-hot agent IDs, MLP hypernetworks do not guarantee distinct weights for each agent. Additionally, they increase the total number of trainable parameters, necessitating a careful balance between expressiveness and computational overhead.
4.2 Agent Embeddings and Initialisation
The agent embedding $e^{i}$ is a one-hot encoded ID for Linear Hypernetworks. For MLP Hypernetworks, we use learned agent embeddings, orthogonally initialised and optimised end-to-end with the hypernetwork. HyperMARL’s hypernetworks are themselves initialised such that the generated agent-specific parameters ($\theta^{i},\phi^{i}$) initially match the distribution of standard direct initialisation schemes (e.g., orthogonal for PPO, preserving fan in/out), promoting stable learning.
4.3 Gradient Decoupling in HyperMARL
A core difficulty in FuPS is cross-agent gradient interference [10, 49]. HyperMARL mitigates this by generating each agent’s parameters through a shared hypernetwork, thereby decoupling agent-conditioned and observation-conditioned components of the gradient.
Hypernetwork gradients. Consider a fully cooperative MARL setting with a centralised critic, we can formulate the policy gradient for agent $i$ as follows [2, 24]:
$\nabla_{\theta^{i}}J(\theta^{i})\;=\;\mathbb{E}_{\mathbf{h_{t}},\mathbf{a_{t}}\sim\boldsymbol{\pi}}\Bigl{[}A(\mathbf{h_{t}},\mathbf{a_{t}})\,\nabla_{\theta^{i}}\log\pi_{\theta^{i}}(a_{t}^{i}\mid h_{t}^{i})\Bigr{]},$
where $\mathbf{h_{t}}$ and $\mathbf{a_{t}}$ are the joint histories and joint actions for all agents, $\theta^{i}$ denotes the parameters of agent $i$, and $A(\mathbf{h_{t}},\mathbf{a_{t}})=Q(\mathbf{h_{t}},\mathbf{a_{t}})-V(\mathbf{h_{t}})$ is the advantage function.
Decoupling. In HyperMARL each agent’s policy weights are produced by the hypernetwork $h_{\psi}^{\pi}$: $\theta^{i}=h_{\psi}^{\pi}(e^{i})$, so we optimise a single parameter vector $\psi$. Applying the chain rule and re-ordering the expectations:
$\nabla_{\!\psi}J(\psi)=\sum_{i=1}^{I}\underbrace{\nabla_{\!\psi}h_{\psi}^{\pi}(e^{i})}_{\mathbf{J}_{i}\;\text{(agent-conditioned)}}\;\underbrace{\mathbb{E}_{\mathbf{h}_{t},\mathbf{a}_{t}\sim\boldsymbol{\pi}}\bigl{[}A(\mathbf{h}_{t},\mathbf{a}_{t})\,\nabla_{\theta^{i}}\log\pi_{\theta^{i}}(a_{t}^{i}\mid h_{t}^{i})\bigr{]}}_{Z_{i}\;\text{(observation-conditioned)}}.$ (4)
  • Agent-conditioned factor $\mathbf{J}_{i}$. This Jacobian depends only on the fixed embedding $e^{i}$ and the hypernetwork weights $\psi$, therefore, it is deterministic with respect to mini-batch samples (as $e^{i}$ and $psi$ are fixed per gradient step), separating agent identity from trajectory noise.
  • Observation-conditioned factor $Z_{i}$. The expectation averages trajectory noise per agent $i$ for its policy component $\pi_{\theta^{i}}$, prior to transformation by $\mathbf{J}_{i}$ and aggregation.
The crucial structural decoupling in Equation (4) ensures HyperMARL first averages noise per agent (via factor $Z_{i}$) before applying the deterministic agent-conditioned transformation $\mathbf{J}_{i}$. This mitigates gradient interference common in FuPS+ID  [10, 49], where observation noise and agent identity become entangled (see Equation (12) in App. F.2). This is the MARL analogue of the task/state decomposition studied by [37, Eq. 18] in Meta-RL. Section 5.2 empirically verifies the predicted variance drop, and ablations in Section 6 demonstrate that disabling decoupling degrades performance, underscoring its critical role.
5 Experiments
Our empirical evaluation of HyperMARL assesses whether agent-conditioned hypernetworks can enable adaptive (specialised or homogeneous) policies without altered RL objectives, preset diversity levels, or sequential updates. We structure our experiments to directly answer two key research questions: Q1: Specialised Policy Learning: Can HyperMARL effectively learn specialised policies via a shared hypernetwork? Q2: Effectiveness in Homogeneous Tasks: Is HyperMARL competitive in environments that necessitate homogeneous behaviours?
To address these questions, HyperMARL is evaluated against five representative modern baselines across a carefully selected suite of MARL benchmarks (18 scenarios, up to 20 agents). All experiments use at least 5 seeds (details in App. G.2.1).
5.1 Experimental Setup
Table 1: MARL environments for evaluating HyperMARL.
Env. Agents Action Behaviour
Dispersion 4 Discrete Hetero.
Navigation 2,4,8 Continuous Homo., Hetero., Mixed
MAMuJoCo 2–17 Continuous Hetero.
SMAX 2–20 Discrete Homo.
Environments. HyperMARL is evaluated across 18 scenarios from four diverse MARL environments (Dispersion [5], Navigation [5], MAMuJoCo [34], and SMAX [36]) (Table 1). These were specifically chosen to rigorously test performance across varying complexities, agent counts (2 to 20), and distinct behaviours (heterogeneous, homogeneous, or mixed). Full details in Appendix G.1.
Baselines. We evaluate HyperMARL against modern parameter sharing (PS) and diversity-promoting baselines. Core PS comparisons use FuPS+ID and NoPS. For specialisation, we include privileged baselines: DiCo [6] (shared and non-shared weights, preset diversity levels); HAPPO [49] (shared critic, sequential actor updates); and Kaleidoscope [26] (learnable masks, critic ensembles, diversity loss). We use IPPO/MAPPO [12, 48] as the underlying algorithm for all methods except Kaleidoscope (MATD3 backbone due to its intricate off-policy design; see App. H.1 for these results).
Adhering to best evaluation practices [33], we  use original codebases/hyperparameters and environments for which baselines were tuned. HyperMARL uses identical observations and generates architectures of equivalent capacity to baselines. Training and evaluation (App. G.2.1) and hyperparameters (App. I) follow each baseline’s original setup. We detail our baseline and environment selection criteria in Table 4, with architecture details in App. G.2.
(a) IPPO
(b) MAPPO
(c) Actor Gradient Var.
(d) Policy Diversity
Figure 5: Performance and gradient analysis. (a,b) IPPO and MAPPO on Dispersion (20M timesteps) - IQM of Mean Episode Return with 95% bootstrap CIs: Hypernetworks match NoPS performance while FuPS struggle with specialisation. Interval estimates in App. H.2.1. (c) Actor gradient variance: Hypernetworks achieve lower gradient variance than FuPS+ID. (d) Policy diversity (SND with Jensen–Shannon distance): Hypernetworks achieve NoPS-level diversity while sharing parameters.
Measuring Policy Diversity. To measure the diversity of the policies we System Neural Diversity (SND) [7] (Equation 5) with Jensen-Shannon distance (details in App. G.2.2).
5.2 Q1: Specialised Policy Learning
Learning Diverse Behaviour (Dispersion) Figures 5(a) and 5(b) show that FuPS variants (IPPO-FuPS, MAPPO-FuPS – ($\bullet$)) can struggle to learn the diverse policies required by Dispersion (even when running for longer - Fig. 19), while their NoPS counterparts (IPPO-NoPS, MAPPO-NoPS–($\bullet$)) converge to the optimal policy, corroborating standard FuPS limitations to learn diverse behaviour. In contrast, HyperMARL (both linear and MLP variants) ($\bullet$, $\bullet$) match NoPS performance, suggesting that a shared hypernetwork can effectively enable agent specialisation. SND policy diversity measurements (Fig. 5(d)) confirm FuPS variants achieve lower behavioural diversity than NoPS, while HyperMARL notably matches NoPS-level diversity.
Figure 6: 17-agent Humanoid learning dynamics (IQM, 95% CI). HyperMARL, utilising a shared actor architecture, outperforms MAPPO-FuPS (non-overlapping CIs) and matches the performance of methods employing non-shared or sequential actors. This challenging environment is recognised for its high variance in outcomes across different methods [49].
Gradient Variance. HyperMARL (IPPO and MAPPO variants) also exhibits lower mean policy gradient variance than FuPS+ID across actor parameters (Fig. 5(c)). This aligns with their ability to learn diverse behaviours and supports the hypothesis that its gradient decoupling mechanism (Sec. 4.3) enhances training stability.
Diversity at Complexity and Scale (MAMuJoCo). In the challenging MAMuJoCo heterogeneous control tasks (Table 2), HyperMARL (MLP variant) is broadly competitive. Notably, unlike HAPPO and MAPPO (independent actors), HyperMARL uses a shared actor and parallel updates, and yet manages strong performance, even in the 17-agent Humanoid-v2 notoriously difficult heterogeneous task[49] (Fig. 6), matching methods that employ independent actors and sequential updates.
Table 2: Mean episode return in MAMuJoCo for MAPPO variants(IQM, 95% CI). HyperMARL achieves the highest IQM in 3/4 scenarios (bold), and is the only method with shared actors to demonstrate stable learning in the notoriously difficult 17-agent Humanoid environment (see Figure 6 for learning dynamics). * indicates CI overlap with the top score.
Scenario HAPPO FuPS+ID Ind. Actors HyperMARL (Ours)
Humanoid-v2 17x1 6501.15* (3015.88, 7229.79) 566.12 (536.36, 603.01) 6188.46* (5006.13, 6851.74) 6544.10 (3868.00, 6664.89)
Walker2d-v2 2x3 4748.06* (4366.94, 6230.81) 4574.39* (4254.21, 5068.32) 4747.05* (3974.76, 6249.58) 5064.86 (4635.10, 5423.42)
HalfCheetah-v2 2x3 6752.40* (6130.42, 7172.98) 6771.21* (6424.94, 7228.65) 6650.31* (5714.68, 7229.61) 7063.72 (6696.30, 7325.36)
Ant-v2 4x2 6031.92* (5924.32, 6149.22) 6148.58 (5988.63, 6223.88) 6046.23* (5924.62, 6216.57) 5940.16* (5485.77, 6280.59)
Adaptability (Navigation). Navigation tasks [5] assess adaptability to homogeneous, heterogeneous, and mixed goals (some agents have the same goals, others different). We compare HyperMARL with baselines including DiCo [6]. While using DiCo’s optimal preset diversity for n=2 agents, we note that identifying appropriate diversity levels for DiCo with larger teams ($n>2$) via hyperparameter sweeps proved challenging (see Tables 12 and 13).
(a) Shared Goals
(b) Unique Goals
(c) Mixed Goals (Half Agents Share Goals)
Figure 7: Average Reward (IQM, 95% CI) in Navigation for IPPO Variants. HyperMARL adapts robustly across goal configurations—(a) shared, (b) unique, and (c) mixed. Both linear and MLP versions consistently match or outperform IPPO baselines and DiCo, with the margin widening as the number of agents grows. Sample-efficiency curves appear in App. H.4.
Across all tested goal configurations (shared, unique, and mixed), HyperMARL consistently achieves strong performance (Figure 7). It generally matches or outperforms NoPS and FuPS, and outperforming DiCo. Interestingly, unlike in sparse-reward tasks like Dispersion, FuPS remains competitive with NoPS and HyperMARL in Navigation scenarios requiring diverse behaviours for smaller teams ($n\in\{2,4\}$), likely due to Navigation’s dense rewards. However, HyperMARL distinguishes itself as the strongest method for n=8 agents, highlighting its effectiveness in handling more complex coordination challenges.
(a) 2s3z
(b) 3s5z
(c) SMACv2 10 Units
(d) SMACv2 20 Units
Figure 8: IQM and 95% CI of mean win rate in SMAX. Performance of FuPS Recurrent IPPO and MAPPO and HyperMARL (MLP) on SMAX. HyperMARL performs comparably to FuPS baselines across all environments, demonstrating its effectiveness in tasks requiring homogeneous behaviours and using recurrency. Inverval estimates in App. 22.
5.3 Q2: Effectiveness in Homogeneous Tasks
SMAX. Finally, we evaluate HyperMARL (MLP) on SMAX, where recurrent FuPS is the established baseline [36, 48, 16]. Figure 8 shows while some FuPS variants might exhibit marginally faster initial convergence on simpler maps, HyperMARL achieves comparable final performance on all maps, using the same GRU backbone for partial observability. These results highlights two points: (i) HyperMARL is fully compatible with recurrent architectures essential under partial observability, and (ii) it has no intrinsic bias toward specialisation and can converge homogenous behaviour when it is optimal (also shown with strong same-goal Navigation performance (Fig. 7(a))), even with large observation spaces and many agents.
Summary. Our empirical results confirm HyperMARL effectively addresses both research questions. For Q1 (Specialisation), across Dispersion, MAMuJoCo, and Navigation, HyperMARL learned specialised policies, matched NoPS-level diversity and performance where FuPS+ID struggled, and scaled to complex, high-agent-count heterogeneous tasks. For Q2 (Homogeneity), HyperMARL demonstrated competitive performance against strong FuPS baselines in SMAX and shared-goal Navigation, confirming its versatility.
6 Ablations
We ablate two critical components of HyperMARL, gradient decoupling(Sec. 4.3) and initialisation scaling (Sec. 4.2) on the complex Humanoid-v2 and simpler Dispersion tasks. Removing gradient decoupling (HyperMARL w/o GD), forcing joint observation/embedding processing, universally degrades performance (Figure 9). Omitting initialisation scaling (HyperMARL w/o RF (reset fan in/out), which aligns initial generated weights with conventional networks, reveals task-dependent effects: it is crucial for the 17-agent Humanoid, echoing prior findings on hypernetwork initialisation [8], but has minor impact in Dispersion. Thus, while principled initialisation becomes vital with increasing complexity, gradient decoupling is fundamentally essential across tested scenarios.
(a) Humanoid w/o GD
(b) Dispersion w/o GD
(c) Humanoid w/o RF
(d) Dispersion w/o RF
Figure 9: Ablation results comparing HyperMARL to variants without gradient decoupling (w/o GD) and without reset fan in/out initialization (w/o RF) across environments. Gradient decoupling (a,b) is consistently critical across both environments, while initialisation scaling (c,d) shows greater importance in the complex Humanoid task but less impact in the simpler Dispersion environment.
7 Related Work
Hypernetworks in RL and MARL. Hypernetworks are effective in single-agent RL for meta-learning, multi-task learning, and continual learning [3, 4, 37, 20]. In MARL, QMIX [35] used hypernetworks (conditioned on a global state) to mix per-agent Q-values; however, each agent’s own network remained a standard GRU. Parallel work, CASH [15], conditions hypernetworks on local observations and team capabilities for zero-shot generalization with heterogeneous action spaces. In contrast, we focus on agent-conditioned hypernetworks for adaptive specialisation within a fixed state-action setting, leveraging gradient decoupling (absent in CASH) that we find critical for specialised behaviours.
Variants of Parameter Sharing. Selective Parameter Sharing (SePS) [10] shares weights between similar groups of agents, identified via trajectory clustering. Pruning methods [23, 26] split a single network into agent-specific subnetworks.
Learning Diverse Policies. Shared parameters often limit policy diversity [10, 23, 16, 25]. Proposed solutions include: (1) maximizing mutual information between agent IDs and trajectories [25], (2) role-based methods [43, 44], or (3) methods that use structural modifications or constraints to induce diversity in agent policies [23, 6, 26]. Outside FuPS/NoPS, HAPPO [49] uses a non-shared centralised critic with individual actors updated sequentially to learn heterogeneous behaviours.
8 Conclusion
We introduced HyperMARL, an approach that uses agent-conditioned hypernetworks to generate per-agent parameters without modifying the standard RL objective or requiring preset diversity levels. Our results show HyperMARL can adaptively learn specialised, homogeneous, or mixed behaviours in settings with up to 20 agents. We also observe a link between HyperMARL’s performance and reduced policy gradient variance, underscoring the importance of decoupling observation- and agent-conditioned gradients. Overall, these findings establish HyperMARL as a promising architecture for diverse MARL tasks. We discuss limitations in App. A, most notably parameter count, which can be remedied by parameter-efficient hypernetworks (e.g., chunked variants [42, 9]).
Appendix A Limitations
Hypernetworks generate weights for target networks, which can lead to high-dimensional outputs and many parameters for deep target networks, particularly when using MLP-based hypernetworks. While HyperMARL uses more parameters than NoPS and FuPS for few agents, it scales almost constantly with agent count, an attractive property for large-scale MARL. Parameter efficiency could be improved through chunking techniques [42, 9], or low-rank weight approximations. This parameter overhead is often acceptable in RL/MARL given typically smaller actor-critic networks, and HyperMARL’s favorable agent scaling (see App. F.3).
Appendix B Broader Impact
This paper presents work whose goal is to advance the field of Multi-Agent Reinforcement Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix C Specialised Policies and Environments
Specialisation plays a key role in MARL, yet remains under-defined, so we define specialised environments and specialised policies.
Definition 1.
An environment is specialised if the following both hold:
  1. 1.
    Distinct Agent Policies. The optimal joint policy $\boldsymbol{\pi}^{*}$ consists of at least two distinct agent policies, i.e., $\exists i,j\in\mathbb{I}$ such that $\pi^{i}\neq\pi^{j}$.
  2. 2.
    Non-Interchangeability. Any permutation $\sigma$ of the policies in $\boldsymbol{\pi}^{*}$, denoted as $\boldsymbol{\pi}^{\sigma}$, results in a weakly lower expected return:
    $\mathbb{E}_{\mathbf{h}\sim\boldsymbol{\pi}^{\sigma}}[G(\mathbf{h})]\leq\mathbb{E}_{\mathbf{h}\sim\boldsymbol{\pi}^{*}}[G(\mathbf{h})],$
    with strict inequality if the joint policies are non-symmetric (i.e., swapping any individual policy degrades performance).
For example, consider a specialised environment such as a football game, optimal team performance typically requires players in distinct roles (e.g., "attackers," "defenders"). Permuting these roles (i.e., exchanging their policies) would typically lead to suboptimal results. Here, agents develop specialised policies by learning distinct, complementary behaviours essential for an optimal joint policy. While agents with heterogeneous capabilities (e.g., different action spaces) are inherently specialised, homogeneous agents can also learn distinct policies. Such environments are analysed in Sections E.1 and 5.2.
Appendix D Measuring Behavioural Diversity
D.1 Quantifying Team Diversity
We quantify policy diversity using System Neural Diversity (SND) [7], which measures behavioural diversity based on differences in policy outputs:
$\operatorname{SND}\left(\left\{\pi^{i}\right\}_{i\in\mathbb{I}}\right)=\frac{2}{n(n-1)|\mathcal{O}|}\sum_{i=1}^{n}\sum_{j=i+1}^{n}\sum_{o\in\mathcal{O}}D\left(\pi^{i}(o),\pi^{j}(o)\right).$ (5)
where $n$ is the number of agents, $\mathcal{O}$ is a set of observations typically collected via policy rollouts, $\pi^{i}(o_{t})$ and $\pi^{j}(o_{t})$ are the outputs of policies $i$ and $j$ for observation $o_{t}$, and $D$ is our distance function between two probability distributions.
In contrast to  [7], we use Jensen-Shannon Distance (JSD) [13, 27] as $D$, rather than the Wasserstein metric [41]. As shown in Appendix D.2, JSD is a robust metric for both continuous and discrete cases, and provides a more reliable measure of policy distance.
D.2 Finding a Suitable Distance Function for Policy Diversity
The choice of distance function $D$ in Equation 5 is crucial for accurately measuring policy diversity. In MARL, policies are often represented as probability distributions over actions, making the choice of distance function non-trivial.
[6] use the Wasserstein metric for continuous policies [41] as distance function $D$, while [29] use the total variation distance for discrete policies. For discrete policies, Wasserstein distance would require a cost function representing the cost of changing from one action to another, which might not be trivial to come up with. On the other hand, although well-suited for discrete policies, TVD might miss changes in action probabilities because it measures the largest difference assigned to an event (i.e. action) between two probability distributions.
We consider a simple example to illustrate this point. Suppose we have two policies $\pi^{1}$ and $\pi^{2}$ with action probabilities as shown in Figure 10. $\pi^{1}$ stays constant, while $\pi^{2}$ changes gradually over timesteps. We see that even as $\pi^{2}$ changes over time, the $TVD(\pi^{1},\pi^{2})$ between $\pi^{1}$ and $\pi^{2}$ remains constant. This is because TVD only measures the largest difference between the two distributions, and does not consider the overall difference between them. On the other hand, the Jensen-Shannon distance (JSD) [13], which is the square root of the Jensen-Shannon divergence, does not have this problem as it is a smooth distance function. Furthermore, it satisfies the conditions for being a metric – it is non-negative, symmetry, and it satisfies the triangle inequality.
For continuous policies, as shown in Figure 11, JSD exhibits similar trends to the Wasserstein distance. Since JSD is a reasonable metric for both continuous and discrete probability distributions, we will use it as the distance metric for all experiments and propose it as a suitable distance function for measuring policy diversity in MARL.
We also summarise the properties of the various distance metrics in Table 3.
Figure 10: Gradual changes in $\pi^{2}$, result in gradual changes in the Jensen-Shannon distance (JSD), while the Total Variation Distance (TVD) can miss changes in action probabilities.
Figure 11: Jensen-Shannon distance (JSD) trends similarly to Wasserstein distance when we have continuous policies.
Method Kinds of Actions Metric Smooth Formula
[…]
Total Variation Distance Discrete Metric No $TV(p,q)=\frac{1}{2}\sum_{x}|p(x)-q(x)|$
Jensen-Shannon Divergence [27] Both Divergence Yes $JSD(p\parallel q)=\frac{1}{2}D_{KL}(p\parallel m)+\frac{1}{2}D_{KL}(q\parallel m),\,m=\frac{1}{2}(p+q)$
Jensen-Shannon Distance [13] Both Metric Yes $\sqrt{JSD(p\parallel q)}$
Table 3: Measure Policy Diversity
Appendix E Specialisation and Synchronisation Games
To study the challenges of specialisation and coordination in an isolated setting, we introduce the Specialisation and Synchronisation Games, drawing inspiration from a version of the XOR game [16], VMAS’s Dispersion [5] and coordination and anti-coordination games in game theory [32]. These environments encourage agents to take distinct actions (Specialisation) or take identical actions (Synchronisation). Despite their deceptively simple payoff structure, these games present substantial learning challenges – non-stationary reward distributions driven by others’ adapting behaviours and in their temporal extension, the joint observation spaces grows exponentially with the number of agents.
E.1 Specialisation and Synchronisation Games Description
Specialisation Game. Agents are encouraged to choose distinct actions. In the simplest setting, it is a two-player matrix game where each agent selects between two actions ($A$ or $B$). As shown in Figure 2(a), agents receive a payoff of $1.0$ when their actions differ (creating two pure Nash equilibria on the anti-diagonal) and $0.5$ when they match. This structure satisfies Definition 1, since optimal joint policies require complementary, non-identical strategies. There is also a symmetric mixed-strategy equilibrium in which each agent plays $A$ and $B$ with probability $0.5$.
Synchronisation Game. Conversely, agents are encouraged to coordinate and choose identical actions. The payoff matrix inverts the Specialisation game’s structure, agents receive $1$ for matching actions and $0.5$ for differing ones. This creates two pure Nash equilibria along the diagonal of the payoff matrix (Figure 2(b)), and incentivises uniform behaviour across agents.
$N$-Agent Extension. Both games naturally scale to $n$ agents and $n$ possible actions. In Specialisation, unique actions receive a payoff of $1.0$, while selecting the same action receives payoffs of $\frac{1}{k}$, where $k$ is the number of agents choosing that action. In contrast, in Synchronisation, agents receive maximum payoffs ($1.0$) only when all actions match. For partial coordination, rewards follow a hyperbolic scale, $\frac{1}{n-k+1}$, encouraging agents to align their choices. Visualisations in Figure 2 and detailed reward profiles appear in Figure 12.
E.2 General-$n$ Payoff Definitions
Both games generalise naturally to $n$ agents and $n$ possible actions. We show the reward profiles for $n=5$ agents in Figure 12.
Let $\mathbf{a}=(a^{1},\dots,a^{n})\in\{1,\dots,n\}^{n}$ and $k_{a}=\bigl{|}\{\,j:a^{j}=a\}\bigr{|}$ be the joint action profile and the count of agents choosing action $a$, respectively.
Temporal Extension. To model sequential decision-making, we extend each normal-form game into a repeated Markov game, where the state at time $t$ is the joint action at time $t-1$. At each step $t$ all agents observe $a^{t-1}$, choose $a_{i}^{t}$, and receive the original Specialisation or Synchronisation payoff. This repeated setup isolates how agents adapt based on past joint actions, exposing temporal patterns of specialisation and coordination.
E.2.1 Specialisation Game
The reward is formulated as follows:
$r_{\mathrm{spec}}^{\,i}(\mathbf{a})=\begin{cases}1.0,&\text{if }k_{a^{\,i}}=1\quad\bigl{(}\text{unique action}\bigr{)};\\[6.0pt] \displaystyle\frac{1}{k_{a^{\,i}}},&\text{if }k_{a^{\,i}}>1\quad\bigl{(}\text{shared action}\bigr{)}.\end{cases}$
E.2.2 Synchronisation Game
The reward is formulated as follows:
$r_{\mathrm{sync}}^{\,i}(\mathbf{a})=\frac{1}{\,n-k_{a^{\,i}}+1\,},$
so that $r_{\mathrm{sync}}^{\,i}=1.0$ when $k_{a^{\,i}}=n$ (all agents select the same action), and otherwise follows a hyperbolic scale encouraging consensus.
Figure 12: Reward profiles for the Specialisation (blue) and Synchronisation (orange) games with $n=5$ agents. In the Specialisation game, an agent’s payoff peaks when it selects a unique action, and then decays as when actions are shared. In the Synchronisation game, payoffs follow a hyperbolic scale $1/(n-k+1)$, reaching maximum only under full consensus, thereby incentivising alignment.
E.3 Proof that FuPS cannot represent the optimal policy in the two-player Specialisation Game
Theorem 1.
A stochastic, shared policy without agent IDs cannot learn the optimal behaviour for the two-player Specialisation Game.
Proof.
Let $\pi$ be a shared policy for both agents, and let $\alpha=\mathbb{P}(a_{i}=0)$ represent the probability of any agent choosing action 0. The expected return of $\pi$ for each agent is:
$E[R(\pi)]$ $=\mathbb{P}(a_{0}=0,a_{1}=0)\cdot 0.5+\mathbb{P}(a_{0}=0,a_{1}=1)\cdot 1$ (6)
$\quad+\mathbb{P}(a_{0}=1,a_{1}=0)\cdot 1+\mathbb{P}(a_{0}=1,a_{1}=1)\cdot 0.5$ (7)
$=0.5\alpha^{2}+2\alpha(1-\alpha)+0.5(1-\alpha)^{2}$ (8)
$=-\alpha^{2}+\alpha+0.5$ (9)
$=-(\alpha-0.5)^{2}+0.75$ (10)
Thus, $E[R(\pi)]\leq 0.75<1$ for all $\alpha\in[0,1]$, with the maximum at $\alpha=0.5$. Therefore, a shared policy cannot achieve the optimal return of 1, confirming the need for specialised behaviour to optimise rewards. ∎
E.4 Measuring Agent Gradient Conflict
We measure gradient conflicts, via the cosine similarity between agents’ gradients $\cos\bigl{(}g_{t}^{(i)},g_{t}^{(j)}\bigr{)}=\frac{\langle g_{t}^{(i)},g_{t}^{(j)}\rangle}{\|g_{t}^{(i)}\|\|g_{t}^{(j)}\|}$, where $g_{t}^{(i)}=\nabla_{\theta}\mathcal{L}^{(i)}(\theta_{t})$.
E.5 Empirical Results in N-player Specialisation and Synchronisation Normal-Form Game
To assess this limitations of FuPS and NoPS in practice, we compare three REINFORCE [46] variants in both games with $n=2,4,8,16,32$ agents: NoPS (No Parameter Sharing), FuPS (Fully Parameter Sharing), and FuPS+ID (FuPS with one-hot agent IDs). All policies use single-layer neural networks with controlled parameter counts (see Appendix I for details).
Figure 13: Mean evaluation reward (mean ± standard error) as a function of the number of agents/actions in the Specialisation (left) and Synchronisation (right) games. In the Specialisation game, vanilla policy gradients (PG, i.e. REINFORCE) with FuPS collapse as the team grows, whereas our identity-aware variant (PG-FuPS+ID) retains near-optimal performance. In the Synchronisation game, PG-NoPS performs well at small scales but degrades with more agents, while both PG-FuPS and PG-FuPS+ID remain at optimal reward across all scales.
Appendix F HyperMARL Details
F.1 HyperMARL Pseudocode
In Algorithm 1, we present the pseudocode for HyperMARL, with HyperMARL-specific steps highlighted in blue. HyperMARL leverages hypernetworks to dynamically generate the parameters of both actor and critic networks. The weights of the hypernetworks and the agent embeddings are automatically updated through automatic differentiation (autograd) based on the computed loss. Additionally, Figure 1 provides a visual representation of the HyperMARL architecture.
Algorithm 1 HyperMARL
1:  Input: Number of agents $n$, number of training iterations $K$, MARL algorithm parameters (e.g., MAPPO-specific hyperparameters)
2:  Initialise:
3:   Hypernetwork parameters $\psi,\varphi$Ensure $\theta^{i}$ and $\phi^{i}$ follow standard initialization schemes, e.g., orthogonal}
4:   Agent embeddings $\{e^{i}\}_{i=1}^{n}$One-hot or orthogonally initialised learnable parameters}
5:  Output: Optimized joint policy $\boldsymbol{\pi}$
6:  for each training iteration $k=0,1,\dots,K-1$ do
7:     for each agent $i=1,\dots,n$ do
8:        $\theta^{i}\leftarrow h_{\psi}^{\pi}(e^{i})$Policy parameters}
9:        $\phi^{i}\leftarrow h_{\varphi}^{V}(e^{i})$Critic parameters}
10:     end for
11:     Interact with environment using $\{\pi_{{\color[rgb]{0.2734375,0.51171875,0.70703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.51171875,0.70703125}\theta^{i}}}\}_{i=1}^{n}$ to collect trajectories $\mathcal{D}$
12:     Compute shared loss $\mathcal{L}$ from $\mathcal{D}$, using $\{V_{{\color[rgb]{0.2734375,0.51171875,0.70703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.51171875,0.70703125}\phi^{i}}}\}_{i=1}^{n}$Standard RL loss function}
13:     Update $\psi$, $\varphi$, and $\boldsymbol{e}$ by minimizing $\mathcal{L}$Optimise parameters.}
14:  end for
15:  Return $\boldsymbol{\pi}=(\pi^{1},\dots,\pi^{n})$
F.2 Variance of the HyperMARL Gradient Estimator
Unbiased estimator.
Following from Equation 4, we can write the unbiased estimator for HyperMARLs gradient as follows:
$\widehat{g}_{\mathrm{HM}}=\sum_{i=1}^{I}\underbrace{\nabla_{\!\psi}h_{\psi}^{\pi}(e^{i})}_{\mathbf{J}_{i}}\;\cdot\;\underbrace{\Bigl{(}\frac{1}{B}\sum_{b=1}^{B}\sum_{t=0}^{T-1}A(\mathbf{h}_{t}^{(b)},\mathbf{a}_{t}^{(b)})\,\nabla_{\theta^{i}}\log\pi_{\theta^{i}}(a_{t}^{i,(b)}\!\mid h_{t}^{i,(b)})\Bigr{)}}_{\widehat{Z}_{i}}=\sum_{i=1}^{I}\mathbf{J}_{i}\,\widehat{Z}_{i},$
where $B$ trajectories $\{\tau^{(b)}\}_{b=1}^{B}$ are sampled i.i.d. and $\widehat{Z}_{i}$ is the empirical analogue of the observation-conditioned factor.
Assumptions.
(A1) trajectories are i.i.d.; (A2) all second moments are finite; (A3) $\psi,\theta,e^{i}$ are fixed during the backward pass.
Variance expansion.
Since each $\mathbf{J}_{i}$ is deterministic under (A3), we may factor them outside the covariance:
$\operatorname{Var}\bigl{(}\widehat{g}_{\text{\tiny HM}}\bigr{)}$ $=\operatorname{Cov}\!\Bigl{(}\sum_{i}\mathbf{J}_{i}\widehat{Z}_{i},\;\sum_{j}\mathbf{J}_{j}\widehat{Z}_{j}\Bigr{)}$ (by def. $\operatorname{Var}(X)=\operatorname{Cov}(X,X)$)
$=\sum_{i,j}\operatorname{Cov}\bigl{(}\mathbf{J}_{i}\widehat{Z}_{i},\;\mathbf{J}_{j}\widehat{Z}_{j}\bigr{)}$ (bilinearity of $\operatorname{Cov}$)
$=\sum_{i,j}\mathbf{J}_{i}\;\operatorname{Cov}(\widehat{Z}_{i},\widehat{Z}_{j})\;\mathbf{J}_{j}^{\!\top}$ (pull deterministic matrices out of $\operatorname{Cov}$)
Equation (LABEL:eq:variance_split_detail) makes explicit that all trajectory-induced covariance is captured $\operatorname{Cov}(\widehat{Z}_{i},\widehat{Z}_{j})$, while the agent-conditioned Jacobians $\mathbf{J}_{i}$ remain trajectory noise-free.
Mini-batch update and covariance.  Let $\widehat{Z}_{i}$ be the unbiased mini-batch estimate of $Z_{i}$ and $\widehat{g}_{\text{\tiny HM}}=\sum_{i}\mathbf{J}_{i}\widehat{Z}_{i}$ the stochastic update. Because every $\mathbf{J}_{i}$ is deterministic (wrt. to mini-batch),
$\operatorname{Var}\!\bigl{(}\widehat{g}_{\text{\tiny HM}}\bigr{)}=\sum_{i,j}\mathbf{J}_{i}\;\operatorname{Cov}(\widehat{Z}_{i},\widehat{Z}_{j})\;\mathbf{J}_{j}^{\!\top},$ (12)
(derivation in Appendix F.2). Equation (12) shows that HyperMARL first averages noise within each agent ($\widehat{Z}_{i}$) and only then applies $\mathbf{J}_{i}$. FuPS+ID, by contrast, updates the shared weights $\theta$ with every raw sample $A\,\nabla_{\theta}\log\pi_{\theta}[h,\mathrm{id}]$, leaving observation noise and agent ID entangled and making it susceptible to gradient interference [10, 49].
F.3 Scalability and Parameter Efficiency
Hypernetworks generate weights for the target network, which can lead to high-dimensional outputs and many parameters for deep target networks. This challenge is amplified in MLP-based hypernetworks, which include additional hidden layers. Figure 14 shows scaling trends:
  • NoPS and linear hypernetworks: Parameter count grows linearly with the number of agents.
  • FuPS: More efficient, as growth depends on one-hot vector size.
  • MLP hypernetworks: Scale better with larger populations, since they only require embeddings of fixed size for each new agent.
To reduce parameter count, techniques like shared hypernetworks, chunked hypernetworks [42, 9], or producing low-rank weight approximations, can be used. While naive implementations are parameter-intensive, this might be less critical in RL and MARL which commonly have smaller actor-critic networks. Moreover, HyperMARL’s near-constant scaling with agents suggests strong potential for large-scale MARL applications.
Figure 14: Parameter scaling for IPPO variants with increasing agents (4 to 1024). MLP Hypernetworks scale nearly constantly, while NoPS, Linear Hypernetworks, and FuPS+One-Hot grow linearly. Log scale on both axes.
Figure 15: Dispersion performance with four agents. FuPS variants match HyperMARL in parameter count but still underperform.
To isolate the effects of parameter count, we scaled the FuPS networks (Figure 15) to match the number of trainable parameters in HyperMARL. Despite generating 10x smaller networks, HyperMARL consistently outperforms FuPS variants, showing its advantages extend beyond parameter count.
F.4 Speed and Memory Usage
We examine the computational efficiency of HyperMARL compared to NoPS and FuPS by measuring inference speed (Figure 16(a)) and GPU memory usage (Figure 16(b)) as we scale the number of agents. The benchmarks were conducted using JAX on a single NVIDIA GPU (T4) with a recurrent (GRU-based) policy architecture. All experiments used fixed network sizes (64-dimensional embeddings and hidden layers) with a batch size of 128 and 64 parallel environments, allowing us to isolate the effects of varying agent count. Each measurement represents the average of 100 forward passes per configuration, with operations repeated across 10 independent trials.
(a) Forward pass inference time
(b) GPU memory usage
Figure 16: Computational efficiency scaling with number of agents. HyperMARL offers a balance between NoPS and FuPS. Notably, in real-world deployments, NoPS incurs additional data transfer and synchronisation costs not reflected here, further widening the efficiency gap.
The results demonstrate that HyperMARL offers a balance between the extremes represented by NoPS and FuPS. In practice, NoPS incurs additional data transfer and update costs, widening the efficiency gap.
Appendix G Experiment Details
G.1 Environments
(a) Dispersion.
(b) Navigation.
(c) MAMuJoCo.
(d) SMAX.
Figure 17: Multi-Agent environments used in our experiments.
Dispersion (VMAS) [5]: A 2D environment where four agents collect unique food particles. This task requires specialised heterogeneous behaviours and resembles the Specialisation Game from Section E.1.
Navigation (VMAS) [5]: Agents navigate in a 2D space to assigned goals, receiving dense rewards based on proximity. Configurations include shared goals (homogeneous), unique goals (heterogeneous), and mixed goals, where some agents share goals while others have unique ones.
Multi-Agent MuJoCo (MAMuJoCo) [34]: A multi-agent extension of MuJoCo, where robot body parts (e.g., a cheetah’s legs) are modelled as different agents. Agents coordinate to perform efficient motion, requiring heterogeneous policies [49].
SMAX (JaxMARL) [36]: Discrete action tasks with 2 to 20 agents on SMACv1- and SMACv2-style maps. FuPS baselines have been shown optimal for these settings [48, 16] indicating homogeneous behaviour is preferred here.
Table 4: Baseline Methods Selection and Justification. Selected methods (✓) were chosen based on their relevance to parameter sharing and specialisation/generalisation in MARL, while excluded methods (✗) did not align with our research objectives or had implementation constraints. Our experimental design systematically addresses key questions on agent specialisation and homogeneity, therefore we selected baselines with demonstrated strong performance in their respective settings, ensuring fair and rigorous comparison.
Method Category Selected Justification & Experimental Settings
IPPO [12]
(NoPS, FuPS+ID) NoPS/FuPS Established MARL baseline implementing both independent (NoPS) and fully shared (FuPS+ID) policy configurations. Tested in: Dispersion, Navigation, SMAX (two SMACv1 maps and two SMACv2 maps, with 10 and 20 agents).
MAPPO [48]
(NoPS, FuPS+ID) NoPS/FuPS Strong baseline with centralized critics for both NoPS and FuPS+ID architectures. Tested in: Dispersion, MAMuJoCo, SMAX (two SMACv1 maps and two SMACv2 maps, with 10 and 20 agents).
DiCo [6] Architectural Diversity Provides comparison with a method employing preset diversity levels that balances shared and non-shared parameters. Tested in: Dispersion and Navigation (as per original paper). Original hyperparameters used for $n=2$ agents; parameter sweeps conducted for $n>2$ to identify optimal diversity levels.
HAPPO [49] Sequential Updates Enables comparison with a method designed for heterogeneous behaviours using sequential policy updates with agent-specific parameters. Tested in: MAMuJoCo, selecting 4/6 scenarios from the original paper, including the challenging 17-agent humanoid task. Walker and Hopper variants were excluded as MAPPO and HAPPO performed similarly in these environments.
Kaleidoscope [26] Architectural Pruning Implemented for off-policy evaluation using its MATD3 implementation with tuned MaMuJoCo hyperparameters. Tested in: MAMuJoCo environments Ant-v2, HalfCheetah-v2, Walker2d-v2 (overlapping with our IPPO experiments), and Swimmer-v2-10x2 (highest agent count variant). Included to evaluate HyperMARL’s competitiveness against a method with ensemble critics and diversity loss, in an off-policy setting.
SEAC [11] Shared Experience Focuses primarily on experience sharing rather than parameter sharing architecture, falling outside our research scope.
SePS [10] Selective Parameter Sharing Requires pretraining phase, which extends beyond the scope of our current study focused on end-to-end learning approaches.
CDAS [25] Intrinsic Reward Only implemented for off-policy methods and has been demonstrated to underperform FuPS/NoPS architectures [16], making it less suitable for our primary on-policy comparisons.
ROMA/RODE [43, 44] Role-based Shows limited practical performance advantages in comparative studies [11], suggesting other baselines provide more rigorous comparison points.
SNP-PS [23] Architectural Pruning No publicly available implementation, preventing direct, reproducible comparison.
G.2 HyperMARL Architecture Details
For the Dispersion and Navigations results (Sec. 5.2) we use feedforward architectures, where we use HyperMARL to generate both the actor and critic networks. For the MAPPO experiments in Section 5.2, for fairness in comparisons with HAPPO and MAPPO, we maintain the centralised critic conditioned on the global state and only use HyperMARL to generate the weights of the actors. For the recurrent IPPO experiments in Section 5.3, HyperMARL only generates the actor and critic feedforward weights, not the GRU weights.
G.2.1 Training and Evaluation
  • Training:
    • For Dispersion (5.2), we run 10 seeds and train for 20 million timesteps.
    • For Navigation (5.2), SMAX (5.3), and MaMuJoCo (5.2), we run 5 seeds and train for 10 million timesteps, consistent with the baselines.
  • Evaluation:
    • For Dispersion (5.2), evaluation is performed every 100k timesteps across 32 episodes.
    • For Navigation (5.2), following the baselines, evaluation is performed every 120k timesteps across 200 episodes.
    • For SMAX (5.3), evaluation is performed every 500k timesteps across 32 episodes.
    • For MaMuJoCo (5.2), following the baselines, evaluation is performed every 25 training episodes over 40 episodes.
G.2.2 Measuring Policy Diversity Details
We measure team diversity using the System Neural Diversity (SND) metric (Equation 5 [7], details Section D) with Jensen-Shannon distance. SND ranges from 0 (identical policies across all agents) to 1 (maximum diversity). We collect a dataset of observations from IPPO-NoPS and IPPO-FuPS policies checkpointed at 5 and 20 million training steps. Each policy is rolled out for 10,000 episodes, generating 16 million observations. We then sample 1 million observations from this dataset to calculate the SND for each method tested.
Appendix H Detailed Results
H.1 Kaleidoscope Off-Policy Comparison Details
Our comparison with Kaleidoscope [26], mentioned in Section 5, is conducted using off-policy methods due to its original design. Kaleidoscope incorporates intricate mechanisms (e.g., learnable masks, an ensemble of five critics, a specific diversity loss) and numerous specialised hyperparameters (e.g., for critic ensembling: ‘critic_deque_len‘, ‘critic_div_coef‘, ‘reset_interval‘; for mask and threshold parameters: ‘n_masks‘, ‘threshold_init_scale‘, ‘threshold_init_bias‘, ‘weighted_masks‘, ‘sparsity_layer_weights‘, etc.). Porting this full complexity to an on-policy PPO backbone would constitute a significant research deviation rather than a direct benchmark of the established method.
Therefore, we utilised Kaleidoscope’s original off-policy implementation to ensure a meaningful comparison. We adopted MATD3 as the algorithmic backbone for this evaluation, as it was the only publicly available Kaleidoscope variant with tuned hyperparameters for Multi-Agent MuJoCo (MaMuJoCo). The MaMuJoCo tasks were chosen for alignment with our primary on-policy (IPPO) results and Kaleidoscope’s original evaluation, specifically: Ant-v2, HalfCheetah-v2, Walker2d-v2 (overlapping with our IPPO experiments), and Swimmer-v2-10x2 ( which represents the MaMuJoCo variant with the highest number of agents). Comparative results in Table 5 show that HyperMARL achieves competitive results with Kaleidoscope, while only using two critics (standard MATD3) and without additional diversity objectives.
Table 5: Mean episode return in MAMuJoCo for off-policy MATD3 variants. IQM of the mean episode returns with 95% stratified bootstrap CI. Bold indicates the highest IQM score; * indicates scores whose confidence intervals overlap with the highest. Although Kaleidoscope employs an ensemble of five critics and an explicit diversity loss, HyperMARL (using a standard MATD3 setup with two critics) achieved competitive results without these additional mechanisms.
Environment Ind. Actors (MATD3) HyperMARL (MATD3) Kaleidoscope (MATD3)
Ant-v2 5270.38 (4329.73, 5719.78) *5886.58 (5840.00, 5920.66) 6160.70 (5798.02, 6463.83)
HalfCheetah-v2 *6777.04 (3169.11, 8233.94) 7057.44 (3508.70, 8818.11) *6901.00 (3609.73, 8192.38)
Walker2d-v2 *5771.87 (5144.84, 8103.34) 7057.68 (5976.50, 8166.09) *6664.32 (5408.95, 8828.11)
Swimmer-v2-10x2 *453.74 (427.24, 487.86) 465.91 (410.82, 475.77) *462.48 (444.22, 475.64)
H.2 Dispersion Detailed Results
H.2.1 Interval Esimates Dispersion
(a) IPPO
(b) MAPPO
Figure 18: Performance of IPPO and MAPPO on Dispersion after 20 million timesteps. We show the Interquartile Mean (IQM) of the Mean Episode Return and the 95% Stratified Bootstrap Confidence Intervals (CI) using [1]. Hypernetworks achieve comparable performance to NoPS, while FuPS struggle with specialisation.
Figure 19: We see that even if we run MAPPO-FuPS on Dispersion for 40 million timesteps (double the timesteps of MLP Hypernetwork), it converges to suboptimal performance.
H.3 Detailed MAMujoco Plots
(a) Humanoid-v2 17x1
(b) Walker2d-v2 2x3
(c) HalfCheetah-v2 2x3
(d) Ant-v2 4x2
Figure 20: Performance of Recurrent IPPO and MAPPO on MaMoJoCo. HyperMARL performs comparably to these baselines, and is the only method with shared actors to demonstrate stable learning in the notoriously difficult 17-agent Humanoid environment.
H.4 Detailed Navigation Plots
(a) 8 Agents, Same Goal
(b) 8 Agents, Different Goals
(c) 8 Agents, Four Goals
(d) 4 Agents, Same Goal
(e) 4 Agents, Different Goals
(f) 4 Agents, Two Goals
(g) 2 Agents, Same Goal
(h) 2 Agents, Different Goals
Figure 21: Sample efficiency of IPPO variants in the VMAS Navigation environment. Plots show IQM and 95% CI (shaded regions) of mean episode return against training steps for different agent counts (rows: 8, 4, 2 agents) and goal configurations (columns, where applicable: Same, Different, Specific Goal Counts). Legend shown at bottom applies to all subplots.
H.5 Interval Esimates - SMAX
(a) 2s3z
(b) 3s5z
(c) SMACv2 10 Units
(d) SMACv2 20 Units
Figure 22: Performance comparison in SMAX environments after 10 million timesteps. We show the Interquartile Mean (IQM) of the Mean Win Rate and the 95% Stratified Bootstrap Confidence Intervals (CI). HyperMARL performs comparably to FuPS baselines across all environments, demonstrating its effectiveness in tasks requiring homogeneous behaviours and using recurrency.
H.6 Additional Ablations
Figure 23: Ablation results comparing HyperMARL with its variants in Dispersion. The results highlight that gradient decoupling is essential for maintaining HyperMARL’s performance.
Appendix I Hyperparameters
Table 6: Hyperparameters, Training and Evaluation for Specialisation and Synchronisation Game
Hyperparameter Value
Environment Configuration
Number of agents 2, 4, 8, 16
Maximum steps per episode 10
Training Protocol
Number of seeds 10
Training steps 10,000
Evaluation episodes 100
Evaluation interval 1,000 steps
Batch size 32
Model Architecture
Hidden layer size 8, 16, 32, 64
Activation function ReLU
Output activation Softmax
Optimization
Learning rate 0.01
Optimizer SGD
Model Parameter Count
2 Agents
NoPS: 60
FuPS: 58
FuPS+ID: 74
FuPS+ID (No State): 42
4 Agents
NoPS: 352
FuPS: 240
FuPS+ID: 404
FuPS+ID (No State): 148
8 Agents
NoPS: 2400
FuPS: 2344
FuPS+ID: 2600
FuPS+ID (No State): 552
16 Agents
NoPS: 17728
FuPS: 17488
FuPS+ID: 18512
FuPS+ID (No State): 2128
Table 7: IPPO and MAPPO Hyperparameters in Dispersion
Hyperparameter Value
LR 0.0005
GAMMA 0.99
VF_COEF 0.5
CLIP_EPS 0.2
ENT_COEF 0.01
NUM_ENVS 16
NUM_STEPS 128
GAE_LAMBDA 0.95
NUM_UPDATES 9765
EVAL_EPISODES 32
EVAL_INTERVAL 100000
MAX_GRAD_NORM 0.5
UPDATE_EPOCHS 4
NUM_MINIBATCHES 2
TOTAL_TIMESTEPS 20000000
ANNEAL_LR false
ACTOR_LAYERS [64, 64]
CRITIC_LAYERS [64, 64]
ACTIVATION relu
SEEDS 30,1,42,72858,2300658
ACTION_SPACE_TYPE discrete
Table 8: MLP Hypernet Hyperparameters in Dispersion
Parameter IPPO MAPPO
HYPERNET_EMBEDDING_DIM 4 8
EMBEDDING_DIM Sweep [4, 16, 64] [4, 8, 16, 64]
HYPERNET_HIDDEN_DIMS 64 64
Table 9: Dispersion Settings
Setting Value
n_food 4
n_agents 4
max_steps 100
food_radius 0.08
share_reward false
penalise_by_time true
continuous_actions false
Table 10: IPPO Hyperparameters for Navigation
Hyperparameters Value
LR 0.00005
NUM_ENVS 600
NUM_STEPS 100
TOTAL_TIMESTEPS $10^{6}$
UPDATE_EPOCHS 45
NUM_MINIBATCHES 30
GAMMA 0.9
GAE_LAMBDA 0.9
CLIP_EPS 0.2
ENT_COEF 0.0
VF_COEF 1.0
MAX_GRAD_NORM 5
ACTIVATION tanh
ANNEAL_LR False
ACTOR_LAYERS [256, 256]
CRITIC_LAYERS [256, 256]
ACTION_SPACE_TYPE continuous
Table 11: MLP Hypernet Hyperparameters in Navigation
Parameter IPPO MAPPO
HYPERNET_EMBEDDING_DIM 4 8
EMBEDDING_DIM Sweep [4, 16, 64] [4, 8, 16, 64]
HYPERNET_HIDDEN_DIMS 64 64
Table 12: DiCo Algorithm $SND\_des$ Hyperparameter
Goal Configuration Number of Agents SND_des
All agents same goal 2 0
4 0
8 0
All agents different goals 2 1.2 (From DiCo paper)
4 [-1,1.2,2.4] $\Rightarrow$ -1 (Best)
8 [-1,1.2,4.8] $\Rightarrow$ -1 (Best)
Some agents share goals 4 [-1,1.2] $\Rightarrow$ -1 (Best)
8 [-1,2.4,1.2] $\Rightarrow$ -1 (Best)
Table 13: Parameter Sweeps for IPPO Variants in Navigation Tasks with Four and Eight Agents
Parameter Sweeps
CLIP_EPS 0.2, 0.1
LR 5e-5, 5e-4, 2.5e-4
Algorithm Setting Selected Values
IPPO-FuPS 8 Agents (Same Goals) 0.2, 5e-5
8 Agents (Different Goals) 0.1, 5e-5
8 Agents (Four Goals) 0.1, 5e-5
4 Agents (Same Goals) 0.2, 5e-5
4 Agents (Different Goals) 0.2, 5e-5
4 Agents (Two Goals) 0.2, 5e-5
IPPO-Linear Hypernetwork 8 Agents (Same Goals) 0.2, 5e-5
8 Agents (Different Goals) 0.1, 5e-5
8 Agents (Four Goals) 0.1, 5e-5
4 Agents (Same Goals) 0.2, 5e-5
4 Agents (Different Goals) 0.1, 5e-5
4 Agents (Two Goals) 0.1, 5e-5
IPPO-MLP Hypernetwork 8 Agents (Same Goals) 0.2, 5e-5
8 Agents (Different Goals) 0.1, 5e-5
8 Agents (Four Goals) 0.1, 5e-5
4 Agents (Same Goals) 0.1, 5e-5
4 Agents (Different Goals) 0.1, 5e-5
4 Agents (Two Goals) 0.1, 5e-5
IPPO-NoPS 8 Agents (Same Goals) 0.1, 5e-5
8 Agents (Different Goals) 0.2, 5e-5
8 Agents (Four Goals) 0.1, 5e-5
4 Agents (Same Goals) 0.1, 5e-5
4 Agents (Different Goals) 0.2, 5e-5
4 Agents (Two Goals) 0.1, 5e-5
IPPO-Dico 8 Agents (Same Goals) 0.2, 5e-5
8 Agents (Different Goals) 0.1, 2.5e-4
8 Agents (Four Goals) 0.1, 2.5e-4
4 Agents (Same Goals) 0.2, 5e-5
4 Agents (Different Goals) 0.1, 2.5e-4
4 Agents (Two Goals) 0.1, 5e-4
Table 14: Environment Settings for Navigation Task
Parameter Value
n_agents 2,4,8
agents_with_same_goal 1, n_agents/2, n_agents
max_steps 100
collisions False
split_goals False
observe_all_goals True
shared_rew False
lidar_range 0.35
agent_radius 0.1
continuous_actions True
Table 15: Default algorithm and model hyperparameters for the Ant-v2-4x2 environment (from  [49]).
Parameter Value
— Algorithm Parameters —
action_aggregation prod
actor_num_mini_batch 1
clip_param 0.1
critic_epoch 5
critic_num_mini_batch 1
entropy_coef 0
fixed_order true
gae_lambda 0.95
gamma 0.99
huber_delta 10.0
max_grad_norm 10.0
ppo_epoch 5
share_param false
use_clipped_value_loss true
use_gae true
use_huber_loss true
use_max_grad_norm true
use_policy_active_masks true
value_loss_coef 1
— Model Parameters —
activation_func relu
critic_lr 0.0005
data_chunk_length 10
gain 0.01
hidden_sizes [128, 128, 128]
initialization_method orthogonal_
lr 0.0005
opti_eps 1e-05
recurrent_n 1
std_x_coef 1
std_y_coef 0.5
use_feature_normalization true
use_naive_recurrent_policy false
use_recurrent_policy false
weight_decay 0
Table 16: Default algorithm and model hyperparameters for the Humanoid-v2-17x1 environment (from  [49]).
Parameter Value
— Algorithm Parameters —
action_aggregation prod
actor_num_mini_batch 1
clip_param 0.1
critic_epoch 5
critic_num_mini_batch 1
entropy_coef 0
fixed_order false
gae_lambda 0.95
gamma 0.99
huber_delta 10.0
max_grad_norm 10.0
ppo_epoch 5
share_param false
use_clipped_value_loss true
use_gae true
use_huber_loss true
use_max_grad_norm true
use_policy_active_masks true
value_loss_coef 1
— Model Parameters —
activation_func relu
critic_lr 0.0005
data_chunk_length 10
gain 0.01
hidden_sizes [128, 128, 128]
initialization_method orthogonal_
lr 0.0005
opti_eps 1e-05
recurrent_n 1
std_x_coef 1
std_y_coef 0.5
use_feature_normalization true
use_naive_recurrent_policy false
use_recurrent_policy false
weight_decay 0
Table 17: Default algorithm and model hyperparameters for the Walker2d-v2-2x3 environment (from  [49]).
Parameter Value
— Algorithm Parameters —
action_aggregation prod
actor_num_mini_batch 2
clip_param 0.05
critic_epoch 5
critic_num_mini_batch 2
entropy_coef 0
fixed_order false
gae_lambda 0.95
gamma 0.99
huber_delta 10.0
max_grad_norm 10.0
ppo_epoch 5
share_param false
use_clipped_value_loss true
use_gae true
use_huber_loss true
use_max_grad_norm true
use_policy_active_masks true
value_loss_coef 1
— Model Parameters —
activation_func relu
critic_lr 0.001
data_chunk_length 10
gain 0.01
hidden_sizes 128, 128, 128
initialization_method orthogonal_
lr 0.001
opti_eps 1e-05
recurrent_n 1
std_x_coef 1
std_y_coef 0.5
use_feature_normalization true
use_naive_recurrent_policy false
use_recurrent_policy false
weight_decay 0
Table 18: Default algorithm and model hyperparameters for the HalfCheetah-v2-2x3 environment (from  [49]).
Parameter Value
— Algorithm Parameters —
action_aggregation prod
actor_num_mini_batch 1
clip_param 0.05
critic_epoch 15
critic_num_mini_batch 1
entropy_coef 0.01
fixed_order false
gae_lambda 0.95
gamma 0.99
huber_delta 10.0
max_grad_norm 10.0
ppo_epoch 15
share_param false
use_clipped_value_loss true
use_gae true
use_huber_loss true
use_max_grad_norm true
use_policy_active_masks true
value_loss_coef 1
— Model Parameters —
activation_func relu
critic_lr 0.0005
data_chunk_length 10
gain 0.01
hidden_sizes 128, 128, 128
initialization_method orthogonal_
lr 0.0005
opti_eps 1e-05
recurrent_n 1
std_x_coef 1
std_y_coef 0.5
use_feature_normalization true
use_naive_recurrent_policy false
use_recurrent_policy false
weight_decay 0
Table 19: HyperMARL Hyperparameters Across MaMuJoCo Environments
Parameter Humanoid Walker2d HalfCheetah Ant Sweeps
v2-17x1 v2-2x3 v2-2x3 v2-4x2
AGENT_ID_DIM 64 64 64 64 None
HNET_HIDDEN_DIMS 64 64 64 64 None
clip_param 0.075 0.0375 0.0375 0.075 [0.1,0.075,0.05,0.0375]
Table 20: Recurrent IPPO and MAPPO Hyperparameters in SMAX (from JaxMARL paper)
Hyperparameter IPPO Value MAPPO Value
LR 0.004 0.002
NUM_ENVS 128 128
NUM_STEPS 128 128
GRU_HIDDEN_DIM 128 128
FC_DIM_SIZE 128 128
TOTAL_TIMESTEPS 1e7 1e7
UPDATE_EPOCHS 4 4
NUM_MINIBATCHES 4 4
GAMMA 0.99 0.99
GAE_LAMBDA 0.95 0.95
CLIP_EPS 0.05 0.2
SCALE_CLIP_EPS False False
ENT_COEF 0.01 0.0
VF_COEF 0.5 0.5
MAX_GRAD_NORM 0.25 0.25
ACTIVATION relu relu
SEED 30,1,42,72858,2300658 30,1,42,72858,2300658
ANNEAL_LR True True
OBS_WITH_AGENT_ID - True
Table 21: Hyperparameter Sweeps and Final Values for Different Maps in SMAX. H- refers to HyperMARL MLP Hypernetworks.
Map Algorithm LR Range Chosen LR HNET Embedding Dim HNET Hidden Dims
2s3z IPPO 0.004 0.004
MAPPO 0.002 0.002
H-IPPO 0.004 0.004 4 32
H-MAPPO 0.002 0.002 64 16
3s5z IPPO 0.004 0.004
MAPPO 0.002, 0.005, 0.0003 0.002
H-IPPO 0.004 0.004 64 16
H-MAPPO 0.002, 0.005, 0.0003 0.0003 64 16
smacv2_10_units IPPO 0.005, 0.001, 0.0003, 0.004 0.001
MAPPO 0.002, 0.005, 0.0003 0.0003
H-IPPO 0.005, 0.001, 0.0003, 0.004 0.005 4 64
H-MAPPO 0.002, 0.005, 0.0003 0.0003 64 16
smacv2_20_units IPPO 0.002, 0.005, 0.0003 0.005
MAPPO 0.002, 0.005, 0.0003 0.0003
H-IPPO 0.002, 0.005, 0.0003 0.005 64 64
H-MAPPO 0.002, 0.005, 0.0003 0.0003 4 64
Note: HNET Embedding Dim refers to the hypernetwork embedding dimension, chosen from the range {4, 16, 64}. HNET Hidden Dims refers to the hidden layer dimensions of the hypernetwork, chosen from the range {16, 32, 64}.
Appendix J Computational Resources
Table 22: Computational Resources by Experiment Type
Experiment Category Hardware Configuration Execution Time Total Hours
Specialisation, 8 cores on AMD EPYC 7H12 2-6 hours per run 250
Synchronisation & Dispersion 64-Core Processor (agent-count dependent)
Navigation 8 cores on AMD EPYC 7H12 4-10 hours per run 320
Experiments 64-Core Processor + NVIDIA RTX A4500
MaMuJoCo 8 cores on AMD EPYC 7H12 8-24 hours per run 1,680
Experiments 64-Core Processor + NVIDIA RTX A4500 (scenario & algorithm dependent)
SMAX 8 cores on AMD EPYC 7H12 2-5 hours per run 160
Experiments 64-Core Processor + NVIDIA RTX A4500